### v2

# TALC: Time-Aligned Captions for Multi-Scene Text-to-Video Generation

Abstract
Recent advances in diffusion-based generative modeling have led to the development of text-to-video (T2V) models that can generate high-quality videos conditioned on a text prompt. Most of these T2V models often produce single-scene video clips that depict an entity performing a particular action (e.g., ‘a red panda climbing a tree’). However, it is pertinent to generate multi-scene videos since they are ubiquitous in the real-world (e.g., ‘a red panda climbing a tree’ followed by ‘the red panda sleeps on the top of the tree’). To generate multi-scene videos from the pretrained T2V model, we introduce Time-Aligned Captions (TALC) framework. Specifically, we enhance the text-conditioning mechanism in the T2V architecture to recognize the temporal alignment between the video scenes and scene descriptions. For instance, we condition the visual features of the earlier and later scenes of the generated video with the representations of the first scene description (e.g., ‘a red panda climbing a tree’) and second scene description (e.g., ‘the red panda sleeps on the top of the tree’), respectively. As a result, we show that the T2V model can generate multi-scene videos that adhere to the multi-scene text descriptions and be visually consistent (e.g., entity and background). Further, we finetune the pretrained T2V model with multi-scene video-text data using the TALC framework. We show that the TALC-finetuned model outperforms the baseline methods by 15.5 points in the overall score, which averages visual consistency and text adherence using human evaluation. The project website is https://talc-mst2v.github.io/.
1 Introduction
The ability to generate videos that simulate the physical world has been a long-standing goal of artificial intelligence [1, 2, 3, 4]. In this regard, text-to-video (T2V) models have seen rapid advancements by pretraining on internet-scale datasets of images, videos, and texts [5, 6]. Previous works [7, 8, 9, 10, 11, 12] primarily focus on training conditional denoising diffusion probabilistic models [13] on paired video-text data [14, 15]. After training, these models allow for video generation by sampling from the trained diffusion model, conditioned on a text prompt. However, most of the open-models such as ModelScope[10] VideoCrafter [16, 17], OpenSora [18] are trained with single-scene video-text dataset [14, 19], which is widely available and easy to acquire. However, real-world scenarios often require the generation of multi-scene videos from multi-scene descriptions (e.g., Scene1: ‘A koala is napping on a tree.’ Scene2: ‘The koala eats leaves on the tree.’). In such cases, the generated video should accurately depict the events in their temporal order (e.g., Scene2 follows Scene1) while maintaining visual consistency, meaning that backgrounds and entities should remain consistent across scenes. While high-performance text-to-video models such as Sora [4] might be able to generate multi-scene videos, we point out that they are closed-source models trained with massive compute resources and lack sufficient details on the model design, training protocol, and datasets. In this work, we present a complementary approach and tackle the challenge of effectively leveraging the capabilities of base T2V models for multi-scene video generation.
The multi-scene text-to-video generation differs from long video synthesis where the goal is to either interpolate (few frames to many frames) [8] or create continuing patterns of the single event in the generated video [11]. Prior works [20, 9] use a transformers [21, 22] to generate video frames for a given scene autoregressively. However, it is hard for their model to generate multiple scenes reliably as the context length increases with history of text descriptions and visual tokens [23] of the previous generated videos (e.g., generating Scene 4 conditioned on the Scene1, 2, 3 videos and descriptions). Other works [24] utilize a latent diffusion model [25] to generate video frames autoregressively by conditioning on the entire history of generated videos and scene descriptions. However, the approach is (a) slow due to repeated sampling, (b) generates only one frame per scene description, and (c) shown to work with only limited cartoon characters [26, 27] instead of wide range of visual concepts in the real-world. In this work, our goal is to generate multi-scene videos in the end-to-end manner, using a diffusion text-to-video generative model that is capable of producing content for a wide range of visual entities and actions.
Figure 1: Multi-scene video generation methods. (a) Generating a video by merging scene 1 and scene 2 descriptions. (b) The resulting video is composed from the video generated by the description of scene 1 and the video generated by the description of scene 2. (c) In our method (TALC) the generated video is conditioned on the description of scene 1 for the first half of the video frames and on the description of scene 2 for the later video frames.
Figure 2: Examples of multi-scene video generation baselines. (a) Generating video on the merged descriptions, leads to a poor text-video alignment. (b) Generating videos for the individual text descriptions and concatenate them temporally, leads to a lack of background consistency. (c) Our approach (TALC) enhances the scene-level text-video alignment and maintains background consistency.
As shown in Figure 1(a), the naive approach to generating a multi-scene video for the scene descriptions $(T^{\prime}_{1},T^{\prime}_{2})$ would condition the T2V generative model on the merged descriptions. In this setup, the diffusion model processes the entire scene description together, and lacks any information regarding the expected temporal order of events in the generated videos. As a result, we find that this approach leads to poor text-video alignment. As shown in Figure 1(b), an alternative approach generates videos for the individual text descriptions independently and concatenates them in the raw input space along the temporal dimension. While this approach achieves good alignment between the scene description and the scene-specific video segment, the resulting video lacks visual consistency in terms of entity and background appearances.
Prior work [28, 29] generates multi-scene videos by utilizing knowledge of the entity, background, and their movements from large language models [30]. However, these videos are generated independently for each scene before being merged. Moreover, these methods do not offer a way to learn from real-world multi-scene video-text data. To remedy these challenges, we propose TALC (Time-ALigned Captions), a simple and effective framework to generate consistent and faithful multi-scene videos. As shown in Figure 1(c), our approach conditions the T2V generative model with the knowledge of the temporal alignment between the parts of the multi-scene video and multi-scene descriptions. Specifically, TALC conditions the visual representations of earlier video frames on the embeddings of the earlier scene description, and likewise, it conditions the representations of later video frames on the embeddings of the later scene description in the temporal dimension. Additionally, the temporal modules in the T2V diffusion architecture allows information sharing between video frames (the first half and the second half) to maintain visual consistency. Thus, TALC enhances the scene-level text-video alignment while providing all the scene descriptions to the diffusion model at once. Further, our TALC framework can enhance the multi-scene text-to-video generation capabilities with real-world multi-scene data (§3.3).
In our experiments, we assess the visual consistency (background and entity consistency) and multi-scene script adherence of the generated videos from Modelscope [10] and Lumiere [6]. Through our automatic and human evaluation, we find that merging scene descriptions leads to high visual consistency but poor text adherence. On the other hand, we observe that merging videos independently achieves the highest text adherence while the visual consistency is compromised. Interestingly, switching to TALC strikes an effective balance between visual consistency and text adherence, outperforming the baseline methods by $11.1$ points on the overall score. This score represents the average of visual consistency and text adherence scores, as determined by human evaluation. Furthermore, we construct a multi-scene text-video dataset from real-world videos and fine-tune the T2V generative model using TALC. On our human evaluation, the generated videos from the TALC-finetuned model exhibit higher text adherence than the base model in multi-scene scenarios. Specifically, it outperforms the baseline methods by $15.5$ points on the overall score.
2 Preliminaries
In this work, we focus on generating multi-scene videos from scene descriptions using a diffusion-based Text-to-Video (T2V) generative model. The initial step is to equip the generative model with the knowledge of a wide range of visual concepts and actions. This is achieved during the pretraining stage (§2.1). Subsequently, we aim to utilize the base model for multi-scene text-to-video generation task, which we formalize in (§2.3). In §3, we propose our TALC framework and discuss collection of real-world multi-scene text-video data for finetuning the base T2V model.
2.1 Diffusion Models for Text-to-Video Generation
Diffusion models [13, 31] $p_{\theta}(x)$ are a class of generative models that learn data distribution $p_{data}(x)$. Due to their flexible design, we can train their class-conditional versions to learn class-conditional data distributions $p_{data}(x|y)$ where $y$ is the conditioning variable, that can take various forms such as labels from a dataset or text description accompanying in a video [32].
We assume a dataset $\mathcal{S}\subset\mathcal{V}\times\mathcal{T}$ consisting of pairs of $(V_{j},T_{j})$ where $V_{j}\in\mathbb{R}^{L\times 3\times H\times W}$ is a raw video consisting of $3$ RGB channels, $L$ frames, $H$ height, $W$ width, and $T_{j}$ is a text caption. We use $\mathcal{V}$ and $\mathcal{T}$ to denote the domain of videos and text, respectively. The aim of T2V generative modeling is to learn the conditional distribution of the videos conditioned on the text $p_{\mathcal{S}}(V_{j}|T_{j})$. In this work, we consider diffusion-based generative models that learn the data distribution via iterative denoising of the input video $z_{j}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$. Here, $z_{j}$ can either represent the input video in the raw pixel space $V_{j}$ [6] or it can represent the latent representation of the video $z_{j}=\mathcal{E}(V_{j})$ for the latent diffusion models [25] where $\mathcal{E}$ is an encoder network such as VAE [33].
Given $z_{j}$, diffused variable $z_{\tau,j}=\alpha_{\tau}z_{j}+\beta_{\tau}\epsilon$ are constructed where $\epsilon\sim\mathcal{N}(0,I)$ where $\alpha_{\tau}$ and $\beta_{\tau}$ are sampled from the noise scheduler $p_{\tau}$ [34] which define the noise levels the model is trained on. Finally, we train a denoiser network $f_{\theta}$ [35, 36] that inputs the diffused variable $z_{\tau}$ and embeddings of the text caption to predict the target vector $y$ where $y$ can be the original noise $\epsilon$, which minimizes the denoising score matching objective [13]:
$\mathbb{E}_{(V_{j},T_{j})\in S,\tau\sim p_{\tau},\epsilon\sim N(0,I)}\left[||\epsilon-f_{\theta}(\tau,z_{\tau,j},h_{j})||_{2}^{2}\right]$ (1)
where $h_{j}=\mathcal{H}(T_{j})\in\mathbb{R}^{d}$ is the embedding of the text caption $T_{j}$ where $\mathcal{H}$ is the text embedding model [37] and $d$ is the dimension size.
2.2 Text Conditioning Mechanism
To ensure the effective textual controllability of video generation, the structure of the denoiser networks is equipped with a cross-attention mechanism [10, 8]. Specifically, it conditions the visual content $z_{\tau}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ on the text. To do so, we first repeat the text embeddings of the text caption $r_{j}=R(h_{j})\in\mathbb{R}^{L\times d}$ where $R$ is a function that repeats the input text embedding $h_{j}$ for $L$ times in the temporal dimension. Intuitively, the repeat operation represents that the $L$ frames of the video $z_{j}$ are semantically aligned with the textual description $T_{j}$ or its text embedding $r_{j}$. In §3, we will manipulate this operation to make the model architecture aware of the video-text alignment in the multi-scene scenario.
These repeated text embeddings $r_{j}$ are inputs to the spatial attention block as the key and value in the multi-head attention block. The cross-attention enables the intermediate visual features to capture the semantic information that facilitates an alignment between the language and vision embeddings. Formally,
$z^{\prime}_{\tau,j}=CA_{f_{\theta}}(Q=z_{\tau,j};K=r_{j};V=r_{j})$ (2)
where $CA_{f_{\theta}}$ is the cross attention mechanism with $Q,K,V$ as the query, key, and value, respectively, in the spatial blocks of the denoiser network. Additionally, $z^{\prime}_{\tau,j}$ is the intermediate representation that is informed with the visual and textual content of the data. In addition to the spatial blocks, the denoiser network also consists temporal blocks that aggregate features across video frames which are useful for maintaining visual consistency in the generated video.
2.3 Multi-Scene Text-to-Video Generation
In many real-world scenarios, such as movies, stories, and instructional videos [38], a video may depict multiple transitions with the same or changing entities, as well as multiple actions or events. In addition, the different video segments often share contextual information such as the background or location. These videos are considered multi-scene videos. In this work, we aim to generate multi-scene video $X=\{x_{1},x_{2},\ldots,x_{n}\}$ from multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ where $n$ are the number of sentences and each sentence $y_{j}$ is a scene description for scene $j$. Additionally, the index $j$ also defines the temporal order of events in the multi-scene script i.e., we want the events described in the scene $j$ to be depicted earlier than the events described in the scene $k$ where $k>j$. Further, we want the parts of the entire generated video $X$, given by $x_{j}$, to have high video-text semantic alignment with the corresponding scene description $y_{j}$, also referred to as text adherence.
For instance, consider a two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’}. Here, we need the T2V generative model to synthesize the appearance of the red panda (an entity) that remains consistent throughout the generated video, also referred to as entity consistency. In addition, we will expect that the context of the multi-scene video of a forest (a background) to remain consistent, also referred to as background consistency.
Figure 3: The architecture of Time-Aligned Captions (TALC). During the generation process of the video, the initial half of the video frames are conditioned on the embeddings of the description of scene 1 ($r_{y_{1}}$), while the subsequent video frames are conditioned on the embeddings of the description of scene 2 ($r_{y_{2}}$).
3 Method
3.1 TALC: Time-Aligned Captions for Multi-Scene T2V Generation
Most of the existing T2V generative models [10, 16, 6] are trained with large-scale short video-text datasets (10 seconds - 30 seconds) such as WebVid-10M [14]. Here, each instance of the dataset consists of a video and a human-written video description. These videos either lack the depiction of multiple events, or the video descriptions do not cover the broad set of events in the video, instead focusing on the major event shown. As a result, the pretrained T2V generative models only synthesize single video scenes depicting individual events.
We introduce TALC, a novel and effective framework to generate multi-scene videos from diffusion T2V generative models based on the scene descriptions. Our approach focuses on the role of text conditioning mechanism that is widely used in the modern T2V generative models (§2.2). Specifically, we take inspiration from the fact that the parts of the generated video $x_{j}$ should depict the events described in the scene description $y_{j}$. To achieve this, we ensure that the representations for the part of the generated video aggregates language features from the scene description $y_{j}$.
Consider that we want to generate a multi-scene video $X\in\mathbb{R}^{L\times 3\times H\times W}$ from the scene descriptions $y_{j}\in Y$, using a T2V generative model $f_{\theta}$. Furthermore, we assume that individual video segments $x_{j}$ are allocated $L/n$ frames within the entire video $X$. Let $z_{X}=[z_{x_{1}};z_{x_{2}};\ldots;z_{x_{n}}]\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ represent the representation for the entire video $X$, and $z_{x_{j}}\in\mathbb{R}^{(L/n)\times C\times H^{\prime}\times W^{\prime}}$ for the $j^{th}$ part of the video that are concatenated in the temporal dimension. In addition, consider $r_{Y}=\{r_{y_{1}},\ldots,r_{y_{n}}\}$ be the set of text embeddings for the multi-scene description $Y$ and $y_{j}$ be an individual scene description. In the TALC framework, the Eq. 2 is changed to:
$z^{\prime}_{\tau,x_{j}}$ $=CA_{f_{\theta}}(Q=z_{\tau,x_{j}},K=r_{y_{j}},V=r_{y_{j}})$ (3)
$z^{\prime}_{\tau,X}$ $=[z^{\prime}_{x_{1}};z^{\prime}_{x_{2}};\ldots;z^{\prime}_{x_{n}}]$ (4)
Here, $\tau$ represents the timestamp in the diffusion modeling setup, which is applied during training as well as inference. We illustrate the framework in Figure 3. While TALC aims to equip the generative model with the ability to depict all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the denoiser network. By design, our approach can be applied to the pretrained T2V model during inference.
3.2 Baselines
Here, we describe the baseline methods that could be used to generate videos for the multi-scene descriptions from a given diffusion text-to-video generative model.
3.2.1 Merging Captions
In this setup, we create a single caption by merging all the multi-scene descriptions. Specifically, the multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ can be written as a single prompt ‘$\mathcal{P}=y_{1}.\text{Then, }y_{2}.\ldots\text{Then, }y_{n}.$’ For instance, the two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’} will change to $\mathcal{P}=$ ‘A red panda climbs on a bamboo forest. Then, the red panda sleeps peacefully in the treetop.’ Subsequently, we generate a video from the T2V model $f_{\theta}$ by conditioning it on $\mathcal{P}$. While this approach mentions the temporal sequence of the events in a single prompt, the T2V model does not understand the temporal boundaries between the two events. Specifically, the Eq. 2 suggests that the visual features for all the video frames will aggregate information from the entire multi-scene description, at once, without any knowledge about the alignment between the scene description and its expected appearance in the generated video.
3.2.2 Merging Videos
In this setup, we generate videos for each scene description individually and merge them in the raw input space. Formally, the individual scene description $y_{i}$ conditions the T2V model $f_{\theta}$ to generate the parts of the multi-video $x_{i}$. Finally, we stitch the individual videos together to synthesize the entire video $X={x_{1},x_{2},\ldots,x_{n}}$. In this process, the parts of the multi-scene video closely adhere to the scene descriptions, leading to high text fidelity. However, since the generated videos do not have access to all the multi-scene descriptions (e.g., the video for Scene 2 is not informed about Scene 1), the visual consistency across the entire video is quite poor.
3.3 Multi-Scene Video-Text Data Generation
Figure 4: Our approach for generating time-aligned video captions. The process begins with PyScene cuts identifying the boundaries of distinct scenes within a video. Keyframes are then selected from the median of each scene. These frames are processed collectively through the Gemini model to produce multi-image captions that maintain narrative continuity by contextualizing each scene within the video’s overall sequence.
While our approach generates better multi-scene videos, the text adherence capabilities of the pretrained T2V generative model are limited. This is due to the lack of multi-scene video-text data during its pretraining. Unlike single video-text datasets, the multi-scene video-text datasets are not widely available and are hard to curate for model training. This is attributed to the fact that high-quality caption generation requires a lot of human labor which is time-consuming and expensive. Prior work such as ActivityNet [39] has curated human captions for specific video scenes depicting useful actions in long videos. However, the video scenes are either overlapping or have a large temporal gap between them that will be harmful for natural and smooth variations between the generated multi-scene videos. Hence, the absence of high-quality captions for continuous video scenes in the dataset makes unsuitable for T2V generative training.
To this end, we aim to create a real-world multi-scene video-text dataset to allow further training of the pretrained T2V models. Specifically, we leverage the capability of the multimodal foundation model, Gemini-Pro-Vision [40], to generate high-quality synthetic data for enhanced video-text training [41]. Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description from the dataset. Subsequently, we utilize PySceneDetect library  [^1] https://github.com/Breakthrough/PySceneDetect to generate continuous video scenes from $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of scene cuts in the video. A similar approach was used in a prior work [12] to detect scene changes in the video data. Then, we sample the middle video frame $F_{i,j}$ as a representative of the semantic content in the video scene $A_{i,j}$. Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for a single video $A_{i}$ and the entire video caption $B_{i}$ to a large multimodal model [40]. Specifically, the model is prompted to generate high-quality captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$. We provide the prompt provided to the multimodal model in Appendix §A. In Figure 4 we provide an instance for the multi-scene video-text data generation.
Datasets. To construct a multi-scene video-text dataset, we utilize existing dataset that include natural (real) videos and associated high-quality human-written captions that summarize the entire video. Specifically, we choose MSR-VTT [42] and VaTeX [43]. Most of the videos in MSR-VTT are 10-30 seconds long while VaTeX consists 10 seconds long videos. In addition, each video in MSR-VTT and VaTex consists 20 captions and 10 captions, respectively, out of which one is selected at random for multi-scene data generation. As described above, a single video is cut into multiple video segments using Pyscene library. In our experiments, we retain the first four video segments and discard any additional segments if the library generates more than four. Since the accuracy of the multi-scene captioning and the computational demands during finetuning are influenced by the number of scenes, we opt to limit the scene count to four for our experiments. However, future work could employ similar methodologies to scale the number of scenes, given more computing power and advanced multi-scene captioning models. We provide the data statistics for the final multi-scene data in Appendix §G.
4 Evaluation
In this section, we describe the evaluation scheme for videos generated from multi-scene text descriptions. First, we describe the evaluation metrics that we aim to assess in this work (§4.1). Then, we generate multi-scene descriptions for a diverse set of tasks (§4.2). Finally, we present the details for automatic and human evaluation of the generated videos (§4.3).
4.1 Metrics
The ability to assess the quality of the generated multi-scene videos is a challenging task itself. As humans, we can judge the multi-scene videos across diverse perceptual dimensions [44] that the existing automatic methods often fails to capture [45]. Following [28], we focus on the visual consistency of the generated video, text adherence capabilities of the T2V models, and video quality of the video. Here, we present the metrics with the aspects that they intend to assess in the generated video for multi-scene text description.
Visual Consistency. This metric aims to assess the (entity or background) consistency between the frames of the multi-scene videos. Here, the entity consistency aims to test whether the entities in the multi-scene video are consistent across the video frames. For instance, the appearance of an animal should not change without a change described in the text description. In addition, the background consistency aims to test whether the background of the multi-scene video remains consistent across the video frames. For instance, the room should not change without a change description in the text.
Text Adherence. This metric aims to test whether the generated video adheres to the multi-scene text description. For instance, the events and actions described in the text script should be presented in the video accurately, and in the correct temporal order.
In our experiments, we compute the visual consistency and text adherence with the automatic and human evaluators. Further, we compute the overall score, which is the average of the visual consistency and text adherence scores. In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Table 1.
4.2 Task Prompts
Here, we curate a set of task prompts for diverse scenarios, aiming to holistically assess the quality of the generated videos.
Single character in multiple visual contexts (S1). In this scenario, we instruct an LLM, GPT-4, to create a coherent script consisting of four scenes. Each scene features a specific animal character performing diverse activities in every scene. This task assesses the capability of the T2V model to generate consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script. For instance, a generated script could be ‘Scene 1: A red panda is climbing a tree. Scene 2: The red panda eats the leaves on the tree. Scene 3: The red panda lies down on the branch of the tree. Scene 4: The red panda sleeps on the branch’. In total, we generate $100$ prompts in this scenario.
Different characters in a specific visual context (S2). In this scenario, we instruct a language model, GPT-4, to create a coherent script consisting of four scenes. Each scene features different animal characters engaging in the same activity in every scene [20]. This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the appearance of the different characters in the multi-scene text script. For instance, a generated script could be ‘Scene 1: A cat leaps onto countertop. Scene 2: A dog leaps onto the same countertop. Scene 3: A rabbit leaps onto the same countertop. Scene 4: A raccoon leaps onto the same countertop’. In total, we generate $100$ prompts in this scenario.
Multi-scene captions from real videos (S3). Here, we aim to assess the ability of the model to generate multi-scene videos for open-ended prompts that are derived from real-world videos. This task also assesses the ability of the T2V model to generate consistent appearances of the entity and its background while adhering to multi-scene descriptions. Specifically, we use our multi-scene video-text data generation pipeline (§3.3) to create such prompts for the real videos from the test splits of the video-text datasets. For example, a multi-scene text script could be ‘Scene 1: A beauty vlogger introduces her skincare routine. Scene 2: She applies a serum to her face, smoothing it in’. We present a sample of the various task prompts in the Appendix §B. In total, we generate $100$ prompts in this scenario.
4.3 Evaluator
In this work, we devise an automatic evaluation framework and perform human evaluation to assess the quality of the multi-scene generated videos.
Automatic Evaluation. Here, we utilize the capability of a large multimodal model, GPT-4-Vision [46], to reason over multiple image sequences. First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for two-scene video). Then, we prompt the multimodal model with the temporal sequence of video frames from different scenes and the multi-scene text description. Specifically, we instruct the multimodal model to decide the quality of the generated video across various metrics including entity consistency, background consistency, and text adherence. For each metric, the multimodal model assigns one of three possible response $\{yes=1,partial=0.5,no=0\}$. For instance, $yes$ for the entity consistency metric implies that the video frames sampled from the generated video have consistent appearance of the entity described in the multi-scene script. In this work, we do not utilize any existing video-text alignment models [47, 41] for evaluating text adherence as they are trained on single-scene video-text datasets. We present the automatic evaluation prompt in Appendix §C.
Human Evaluation. We also conduct a human evaluation to assess the multi-scene generated videos along the dimensions of visual consistency, text adherence, and visual quality. Specifically, we ask the annotators from Amazon Mechanical Turk (AMT) to choose one of three options for each metric $\{yes,partial,no\}$, similar to the automatic evaluation. In addition, we choose the annotators that pass a preliminary qualification exam. We present the screenshot of the UI in Appendix §D.
4.4 Evaluation Setup
Since merging captions (§3.2) and TALC3.1) methods input the entire multi-scene text description at once, the quality of the video generated by these methods is influenced by the number of scenes described in the text script. Hence, we calculate the performance of the baselines and TALC by averaging the scores assigned to videos generated for two, three, and four scenes. Additionally, we report on visual consistency by averaging the performance across the entity and background consistency metrics. Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), and the background consistency and text adherence scores are computed for all the task prompts. We also evaluate the impact of TALC-based finetuning on the single scene generation in Appendix §I.
5 Experiments
5.1 Text-to-Video Generative Models
In this work, we utilize ModelScope [10] and Lumiere [6] T2V models for multi-scene video generation. Here, ModelScope is an open-source T2V model with 1.7 billion parameters including the video encoder, text encoder, and denoising U-net network. Specifically, it is trained to generate 16 video frames on the mix of WebVid [14] video-text dataset and LAION [48] image-text dataset. We perform most of our experiments on ModelScope due to its easy-of-access and adoption in prior works [28]. In addition, we also include Lumiere-T2V, a model that leverages space-time U-Net denoising networks to generate high-quality videos. In this work, we include early experiments with Lumiere to showcase the flexibility of the TALC approach for multi-scene video generation.
Base model with TALC. As described in §3.1, our approach modifies the traditional text-conditioning mechanism to be aware of the alignment between text descriptions and individual video scenes. By design, the TALC framework can be applied to the base T2V model during inference, without any multi-scene finetuning. Thus, we compare the performance of the multi-scene videos generated from ModelScope and Lumiere T2V base models under three settings: merging captions, merging videos, and TALC. In this setting, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere. We provide more details on the inference in Appendix §F.
Finetuning with TALC. Since the base model is pretrained with single-scene data, we aim to show the usefulness of TALC framework when we have access to the multi-scene video-text data. To this end, we finetune ModelScope on the multi-scene video-text data (§3.3) with TALC framework. As a pertinent baseline, we also finetune the ModelScope without TALC framework by naively merging the scene-specific captions in the raw text space. In this setting, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$. We provide further details on the finetuning setup in Appendix §H. The inference settings are identical to the prior method of generating videos from the base model without finetuning.
(a) Performance on ModelScope T2V model.
(b) Performance on Lumiere T2V model.
Figure 5: Automatic evaluation results for (a) ModelScope and (b) Lumiere. In (a), we observe that TALC-finetuned ModelScope model achieves the highest overall score, that is the average of the visual consistency and text adherence scores. In (b), we find that TALC framework with the Lumiere base model outperforms merging captions and merging videos on the overall scores. We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
In this section, we present the results for the baselines and TALC framework averaged over a diverse task prompts and multiple scenes using automatic evaluation (§5.2) and human evaluation (§5.3). Finally, we provide qualitative examples for the multi-scene generated videos to showcase the usefulness of our approach (§5.4).
5.2 Automatic Evaluation
We compare the performance of the baselines (e.g., merging captions and merging videos) with the TALC framework for ModelScope and Lumiere using the automatic evaluation in Figure 5.
TALC outperforms the baselines without any finetuning.
In Figure 5(a), we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($68.6$ points), outperforms the overall score achieved by the videos generated using merging captions ($61.7$ points) and merging videos ($67.5$ points) with the base ModelScope. Specifically, we observe that the visual consistency of the generated video is high for merging captions ($91$ points) and TALC ($89.9$ points) while it is low for merging videos ($65$ points). This indicates that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames. In addition, we observe that the text adherence using TALC outperforms merging captions by $14.8$ points, while the text adherence is the highest with a score of $70$ points using merging videos. This can be attributed to the design of the merging videos baseline where individual video scenes adhere to the scene-specific descriptions well. Hence, merging videos independently approach can be viewed as an upper bound on the text adherence metric.
In Figure 5(b), we observe similar trends for the Lumiere T2V generative model. Specifically, we find that the overall score for TALC outperforms merging captions and merging videos by $4$ points and $2$ points, respectively. In addition, we observe that merging captions and TALC achieve a high visual consistency score while merging videos independently has poor visual consistency. Further, we find that TALC outperforms merging captions by $5$ points on text adherence, while merging videos achieves the highest text adherence $65$ points. This highlights that the model more easily generates multi-scene videos that adhere to individual text scripts, whereas adherence to the text diminishes when the model is given descriptions of multiple scenes all at once.
Finetuning with TALC achieves the best performance.
Earlier, we evaluated the usefulness of the TALC framework with the base model. However, the base models are trained with the single-scene video-text data that might limit their capability for multi-scene video generation. To alleviate this issue, we finetune ModelScope T2V model on the multi-scene video-text data (§3.3). Specifically, we finetune the model using the merging captions method and TALC framework, independently.
In Figure 5(a), we find that finetuning with TALC achieves the highest overall score of $75.6$ points in comparison to all the baselines. Specifically, we observe that the visual consistency does not change much with finetuning using the TALC method ($89.9$ points vs $89$ points). Interestingly, we observe that finetuning with merging captions reduces the visual consistency by a large margin of $14$ points. This can be attributed to the lack of knowledge about the natural alignment between video scenes and individual scene descriptions, which gets lost during the merging of captions.
Additionally, we find that the text adherence of the TALC-finetuned model is $15.1$ points more than the text adherence of the TALC-base model. Similarly, we find that the text adherence of the merging captions-finetuned model is $5.1$ points more than the text adherence of the merging captions-base model. This highlights that finetuning a T2V model with multi-scene video-text data helps the most with enhancing its text adherence capability.
Fine-grained Results.
To perform fine-grained analysis of the performance, we assess the visual consistency and text adherence scores for the baselines and TALC framework across diverse task prompts and number of scenes on ModelScope. We present their results in Appendix §E. In our analysis, we find that finetuning with TALC achieves the highest overall score over the baselines across all the scenarios. In addition, we notice that the highest performance is achieved in the scenario that consist of the different entities in a specific visual context. Further, we observe that the performance of the all the methods reduces when the task prompts get more complex i.e., multi-scene captions from real videos. In addition, we observe that finetuning with TALC achieves the highest overall score over the baselines across all the number of scenes. Specifically, we observe that the performance of the merging captions and TALC framework reduces as the number of scenes being generated increases. Overall, we show that the TALC strikes a good balance between visual consistency and text adherence to generate high-quality multi-scene videos.
5.3 Human Evaluation
Figure 6: Human evaluation results for ModelScope model. We observe that the base model using the TALC framework outperforms the merging captions and merging videos baselines on the overall score. In addition, TALC-finetuned model enhances the text adherence and achieves the highest overall score. We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
Table 1: Human evaluation results on the visual quality of the generated videos from ModelScope. We observe that the visual quality of the generated videos are close to each other for the base model. However, finetuning the model with merging captions reduces the video quality by a large margin while TALC-finetuned model retains the video quality.
Method Quality
Merging Captions (Base) 80.5
Merging Videos (Base) 86.5
TALC (Base) 84.5
Merging Captions (F.T.) 63.4
TALC (F.T.) 83.3
TALC achieves the best performance in human evaluation.
We compare the performance of the baselines and TALC framework for ModelScope using human evaluation in Figure 6. We find that TALC-finetuned model outperforms the merging captions and merging video methods with the base model by $12$ points and $15.5$ points, respectively, on the overall score. In addition, we find that using TALC framework in the base model outperforms the merging captions and merging video methods with the base model by $7.6$ points and $11.1$ points, respectively, on the overall score. Further, we observe that the merging captions with the base model achieves the highest visual consistency score of $96.5$ points while it is the lowest for merging videos generated from the base model. In addition, we find that the text adherence of the TALC-finetuned and TALC-base model is better than merging captions-finetuned and merging captions-base model, respectively. Our results highlight at the benefit of including the inductive bias of temporal alignment between the video scenes and their scene descriptions for multi-scene video generation.
Visual quality of the generated videos.
We compare the visual quality of the generated videos using human evaluation in Table 1. We find that the visual quality of videos generated from the base model ranges from $80.5-86.5$ using the baselines and TALC framework. However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points. This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video. Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining. While our work is centered around multi-scene evaluation, we also perform single-scene evaluation in Appendix §I.
5.4 Qualitative Analysis
We provide qualitative examples of generating multi-scene videos using ModelScope (fine-tuned with TALC) and Lumiere (base model with TALC) for diverse scenarios in Figure 12. Our analysis reveals that both ModelScope and Lumiere are capable of producing multi-scene videos that exhibit high text adherence and visual consistency.
Considering the case of the same animal engaging in multiple actions (referred to as "one character multiple contexts"). The videos generated by ModelScope successfully maintained the same animal while varying the background and action between the scenes. Conversely, the videos generated by Lumiere displayed the same animal performing different actions with minimal background alterations. We believe that this distinction is attributed to ModelScope’s fine-tuning with TALC. Considering different animals within a particular visual setting (referred to as "multiple-characters same context"), both ModelScope and Lumiere demonstrated impressive abilities in preserving the consistency of the background across the videos and adhering closely to the provided text. During our analysis, we noticed that the multi-scene captions derived from real videos (referred to as "open-ended captions") exhibited a substantial number of changes between the various scenes. In this scenario, Lumiere, when employed without fine-tuning, displayed challenges in adhering to the text, while ModelScope achieved a higher degree of text adherence but was also prone to visual artifacts.
6 Related Work
Text-to-Video Generative Modeling.
The field of text-to-video (T2V) synthesis has significantly evolved from its inception with models like VGAN [2] and MoCoGAN [49], leveraging the foundational technologies of GANs [50] and VAEs [51] to produce concise, single-scene videos. The narrative depth was further expanded through transformer-based architectures such as CogVideo [52] and VideoGPT [53], enhancing the complexity of video content yet remaining within the confines of single scenes. The advent of diffusion models, exemplified by Imagen Video [54], marked a notable advancement in T2V synthesis. Despite these strides, the challenge of creating multi-scene videos that reflect the complexity of the physical world [1, 2, 3] remains. Our work, TALC, extends the capabilities of T2V models to multi-scene storytelling, filling a crucial gap in the synthesis landscape.
Image-to-Video Animation.
The exploration of multi-scene video generation, innovative methods such as Lumiere [6] and Make-a-Video [55] have employed a two-step process, transforming text to images and then animating these images into videos. While these approaches have advanced visual quality, they often fall short in weaving seamless multi-scene narratives. This limitation is echoed in the work of Emu Video [8], which underscores the difficulty of achieving narrative coherence across multiple scenes. TALC focuses on direct generation of multi-scene narratives from textual prompts aiming for a narrative flow and visual consistency across scenes.
Multi-Scene Video Generation.
The pursuit of multi-scene T2V synthesis has been furthered by recent innovations like Phenaki [20] and Stable Video Diffusion [12], which have explored new frontiers in video generation from textual prompts and the scaling of latent diffusion models, respectively. Additionally, Dreamix [56] and Pix2Video [57] have broadened the scope of diffusion models, applying them to video editing and animation. Despite these advancements, the task of generating videos that convey coherent narratives across multiple scenes remains formidable, highlighted by recent works such as VideoPoet [9], ModelScope [10] and Make-A-Scene [58]. TALC tackles this task and offers a framework produces videos spanning multiple scenes. We also introduce nuanced evaluation approach. This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis.
7 Conclusion
We introduced TALC, a simple and effective method for improving the text-to-video (T2V) models for multi-scene generation. Specifically, it incorporates the knowledge of the natural alignment between the video segments and the scene-specific descriptions. Further, we show that TALC-finetuned T2V model achieve high visual consistency and text adherence while the baselines suffer from one or both of the metrics. Given its design, our framework can be easily adapted into any diffusion-based T2V model. An important future direction will be to scale the amount of multi-scene video-text data and deploy TALC framework during pretraining of the T2V models.
8 Acknowledgement
We would like to thank Ashima Suvarna for providing feedback on the draft. Hritik Bansal is supported in part by AFOSR MURI grant FA9550-22-1-0380.
Appendix A Prompt for Multi-Scene Caption Generation
We present the prompt used to generate multi-scene captions using large multimodal model, Gemini-Pro-Vision, in Figure 7. In particular, we utilize Gemini-Pro-Vision since it can reason over multiple image sequences. Specifically, we provide the multimodal model with a video frame from each of the segmented videos and the single caption for the entire video present in the original video-text datasets.
Your task is to create captions for a series of images, each taken from different video scenes. For every image shown, craft a 7-10 word caption (7-10 words) that precisely describes what’s visible, while also linking these captions into a fluid, engaging story. A common caption will be given to help guide your narrative, ensuring a smooth transition between scenes for a cohesive story flow. Remember to not hallucinate in your responses.
Common Caption: {caption}
Figure 7: Prompt to generate multi-scene caption using large multimodal models.
Appendix B Task Prompts
B.1 Single character in multiple visual contexts
We present the prompt used to generate multi-scene text descriptions for single character in multiple visual contexts from GPT-4 in Figure 8.
Create four concise continuous movie scenes (7-10 words) focusing on a specific real-world character. The scenes should form a cohesive narrative.
Guidelines:
Choice of Character: Select a real-world animal as the focal point of your scenes.
Scene Description: Clearly describe the setting, actions, and any notable elements in each scene.
Connection: Ensure that the scenes are logically connected, with the second scene following on from the first.
Brevity and Precision: Keep descriptions short yet vividly detailed.
Example:
Character: polar bear
Scene 1: A polar bear navigates through a icy landscape.
Scene 2: The polar bear hunts seals near a crack in the ice.
Scene 3: The polar bear feasts on the seal.
Scene 4: The polar bear curls up for a nap.
Now it’s your turn.
Figure 8: GPT-4 Prompt to generate multi-scene prompts for single character in multiple visual contexts.
B.2 Different characters in a specific visual context
We present the prompt used to generate multi-scene text descriptions for different characters in a specific visual context from GPT-4 in Figure 9.
Create four concise scene descriptions (7-10 words) where different characters perform identical action/events.
Choice of Characters: Select four real-world animals as the focal point of the individual scenes.
Background Consistency: Ensure that the background is consistent in both the scenes.
Brevity and Precision: Keep descriptions short yet vividly detailed.
Example:
Characters: teddy bear, panda, grizzly bear, polar bear
Scene 1: A teddy bear swims under water.
Scene 2: A panda swims under the same water.
Scene 3: A panda swims under the same water.
Scene 4: A panda swims under the same water.
Now it’s your turn.
Figure 9: GPT-4 Prompt to generate multi-scene prompts for different characters in a specific visual context.
Appendix C Automatic Evaluation Prompt
We present the prompt used to perform automatic evaluation of the multi-scene generated videos using large multimodal model, GPT-4-Vision, in Figure 10. We utilize GPT-4-Vision for automatic evaluation since it can reason over multiple images. Specifically, we provide the multimodal model with four video frames for each scene in the generated video. The model has to provide its judgments based on the entity consistency, background consistency, and text adherence of the video frames.
You are a capable video evaluator. You will be shown a text script with two-scene descriptions where the events/actions . Video generating AI models receive this text script as input and asked to generate relevant videos. You will be provided with eight video frames from the generated video. Your task is to answer the following questions for the generated video.
1. Entity Consistency: Throughout the video, are entities consistent? (e.g., clothes do not change without a change described in the text script)
2. Background Consistency: Throughout the video, is the background consistent? (e.g., the room does not change described in the text script)
3. Text Adherence: Does the video adhere to the script? (e.g., are events/actions described in the script shown in the video accurately and in the correct temporal order)
Respond with NO, PARTIALLY, and YES for each category at the end. Do not provide any additional explanations.
Two-scene descriptions:
Scene 1: {scene1}
Scene 2: {scene2}
Figure 10: Prompt used to perform automatic evaluation of the multi-scene generated videos. We use this prompt when the number of scenes in the task prompt is two.
Appendix D Human Evaluation Screenshot
We present the screenshot for the human evaluation in Figure 11. Specifically, we ask the annotators to judge the visual quality, entity consistency, background consistency, and text adherence of the multi-scene generated videos across diverse task prompts and number of scenes.
Figure 11: Human Annotation Layout
Appendix E Fine-grained Analysis
We present the automatic evaluation results across task prompts (§E.1) and number of scenes (§E.2).
E.1 Task Prompts
We compare the performance of the baselines and TALC framework across different task prompts in Table 2. We find that the TALC-finetuned model achieves the highest overall score over all the baselines. Specifically, we find that the TALC framework achieves a high visual consistency with scores close to the merging captions baseline. Further, we observe that the TALC framework achieves a higher text adherence in comparison to the merging captions, with or without finetuning, across all the task prompts.
Table 2: Automatic evaluation results across task prompts. Here, S1 refers to the single character in multiple visual contexts. S2 refers to the different characters in a specific visual context. S3 refers to the multi-scene captions from real videos. We abbreviate Finetuning as F.T., Visual consistency as V.C., Text adherence as T.A.
Method S1 S2 S3
V.C. T.A. Overall V.C. T.A. Overall V.C. T.A. Overall
Merging captions (Base) 95.9 43.8 69.9 99.5 21.5 60.5 81.9 32.0 56.9
Merging videos (Base) 69.4 71.5 70.5 71.2 82.3 76.7 57.9 58.7 58.3
TALC (Base) 93.6 48.3 71.0 99.3 57.3 78.3 81.4 36.1 58.8
Merging captions (F.T.) 94.2 57.1 75.7 94.9 31.2 63.1 50.8 24.3 37.5
TALC (F.T.) 93.3 62.6 77.9 98.9 76.3 87.6 79.7 47.9 63.8
E.2 Number of Scenes
We compare the performance of the baselines and TALC framework across different number of generated scenes in Table 3. We find that the TALC-finetuned model outperforms all the baselines in this setup. In addition, we find that the visual consistency of the TALC framework does not change much with the number of the scenes. However, we find that the text adherence of the baselines and TALC framework reduces with the number of generated scenes. The text adherence scores of the merging videos does not change with the number of scenes as it generates the videos for the individual scenes independently.
Table 3: Automatic evaluation results across different number of scenes in the task prompts. We abbreviate Finetuning as F.T., visual consistency as V.C., and Text Adherence as T.A.
# scenes = 2 # scenes = 3 # scenes = 4
V.C. T.A. Overall V.C. T.A. Overall V.C. T.A. Overall
Merging captions (Base) 93.2 34.4 63.8 92.5 33.0 62.7 87.4 29.9 58.6
Merging videos (Base) 66.7 69.9 68.3 65.2 71.9 68.6 63.5 70.7 67.1
TALC (Base) 92.6 54.4 73.5 89.8 48.0 68.9 87.3 39.3 63.3
Merging captions (F.T.) 87.7 45.7 66.7 83.2 39.5 61.4 60.0 27.4 43.7
TALC (F.T.) 88.5 66.6 77.5 90.7 64.1 77.4 87.8 56.1 71.9
Appendix F Inference Details
We provide the details for sampling multi-scene videos from the ModelScope and Lumiere T2V models in Table 4 and Table 5, respectively.
Table 4: Sampling setup for ModelScope T2V model.
Resolution 256 $\times$ 256
Number of video frames per scene 16
Guidance scale 12
Sampling steps 100
Noise scheduler DPMSolverMultiStepScheduler
Table 5: Sampling setup for Lumiere T2V model.
Resolution 1024 $\times$ 1024
Number of video frames per scene 80
Guidance scale 8
Sampling steps 256
Noise scheduler DPMSolverMultiStepScheduler
Appendix G Multi-Scene Data Statistics
We provide the details for the multi-scene video-text dataset in Table 6.
Table 6: Multi Scene Video-Text Data Statistics
Number of entire videos 7107
% of single scene 27.3%
% of two scenes 25.4%
% of three scenes 31.3%
% of four scenes 16.0%
Number of video scene-caption instances 20177
Appendix H Finetuning Details
We provide the details for finetuning ModelScope T2V model with TALC framework in Table 7.
Table 7: Training details for the TALC-finetuned ModelScope T2V model.
Base Model ModelScope  [^2] https://huggingface.co/damo-vilab/text-to-video-ms-1.7b/tree/main
Trainable Module UNet
Frozen Modules Text Encoder, VAE
Batch size 20
Number of GPUs 5 Nvidia A6000
Resolution 256 $\times$ 256
Crop CenterCrop
Learning Rate Scheduler Constant
Peak LR 1.00E-05
Warmup steps 1000
Optimizer Adam (0.9, 0.999, 1e-2, 1e-8) [59]
Max grad norm 1
precision fp16
Noise scheduler DDPM
Number of frames per video scene 8
prediction type epsilon
Appendix I Single-Scene Video Generation
Figure 12: Examples of videos generated by Time-Aligned Captions (TALC)
To ascertain that our model’s new multi-scene generation function does not detract from its single-scene generation performance, we conducted a series of evaluations using the VBench framework [44]. VBench offers a robust analysis of various video generation aspects such as adherence to text prompts, stylistic integrity, semantic coherence, and overall aesthetic quality.
Our analysis, shown in Table 8, establishes a refined baseline: ModelScope (Single-Scene Finetuning), fine-tuned on single-scene video generation data. This process yielded an average score of 0.48, indicating a decrease from the base ModelScope’s average score of 0.63. This suggests that the optimizations in the base model, such as integrating high-quality images, are not fully utilized in single-scene fine-tuning.
Interestingly, fine-tuning the model on multi-scene data (ModelScope - Multi-Scene Finetuning) resulted in improved performance with an average score of 0.59, surpassing the single-scene fine-tuned version. This indicates that multi-scene data enriches the model’s understanding of video content, enhancing both multi-scene and single-scene video generation.
This comparison highlights the importance of data curation and fine-tuning strategies, showing that our approach not only enables complex multi-scene narratives but also improves single-scene video generation.
Table 8: Single-Scene Evaluation Results using VBench, comparing the base model, when it is fine-tuned on multi-scene data, and single-scene data (‘f.t.’ stands for fine-tuned). Our analysis shows ModelScope (Single-Scene Finetuning) as a refined baseline with an average score of 0.48, compared to the base ModelScope’s 0.63. Fine-tuning on multi-scene data (ModelScope - Multi-Scene Finetuning) yields an improved score of 0.59, highlighting the efficacy of multi-scene data in enhancing video generation performance.
Dimension ModelScope (Base) ModelScope (Multi-scene f.t.) ModelScope (Single-Scene f.t.)
Appearance style 0.23 0.24 0.21
Color 0.85 0.83 0.78
Human action 0.96 0.92 0.75
Object class 0.86 0.77 0.42
Overall consistency 0.26 0.26 0.22
Spatial relationship 0.35 0.29 0.14
Subject consistency 0.90 0.83 0.75
Temporal flickering 0.97 0.89 0.86
Temporal style 0.26 0.25 0.21
Average 0.63 0.59 0.48

### v3

@@
 However, it is pertinent to generate multi-scene videos since they are ubiquitous in the real-world (e.g., ‘a red panda climbing a tree’ followed by ‘the red panda sleeps on the top of the tree’).
-To generate multi-scene videos from the pretrained T2V model, we introduce Time-Aligned Captions (TALC) framework.
+To generate multi-scene videos from a pretrained T2V model, we introduce Time-Aligned Captions (TALC) framework.
 Specifically, we enhance the text-conditioning mechanism in the T2V architecture to recognize the temporal alignment between the video scenes and scene descriptions.
-For instance, we condition the visual features of the earlier and later scenes of the generated video with the representations of the first scene description (e.g., ‘a red panda climbing a tree’) and second scene description (e.g., ‘the red panda sleeps on the top of the tree’), respectively.
-As a result, we show that the T2V model can generate multi-scene videos that adhere to the multi-scene text descriptions and be visually consistent (e.g., entity and background).
-Further, we finetune the pretrained T2V model with multi-scene video-text data using the TALC framework.
-We show that the TALC-finetuned model outperforms the baseline methods by 15.5 points in the overall score, which averages visual consistency and text adherence using human evaluation.
-The project website is https://talc-mst2v.github.io/.
+As a result, we show that the pretrained T2V model can generate multi-scene videos that adhere to the multi-scene text descriptions and be visually consistent (e.g., w.r.t entity and background).
+Our TALC-finetuned model outperforms the baseline methods on multi-scene video-text data by 15.5 points on aggregated score, averaging visual consistency and text adherence using human evaluation.
+The project website is https://talc-mst2v.github.io/
+Figure 1: Multi-scene video generation methods.
+(a) Merging scene descriptions; (b) Generating a video per scene independently and concatenating the resulting videos.; (c) TALC: conditioning the generated video on scene-1 description for the first half of the video frames and on scene-2 description for the later video frames.
 1 Introduction
 The ability to generate videos that simulate the physical world has been a long-standing goal of artificial intelligence [1, 2, 3, 4].
-In this regard, text-to-video (T2V) models have seen rapid advancements by pretraining on internet-scale datasets of images, videos, and texts [5, 6].
-Previous works [7, 8, 9, 10, 11, 12] primarily focus on training conditional denoising diffusion probabilistic models [13] on paired video-text data [14, 15].
-After training, these models allow for video generation by sampling from the trained diffusion model, conditioned on a text prompt.
-However, most of the open-models such as ModelScope[10] VideoCrafter [16, 17], OpenSora [18] are trained with single-scene video-text dataset [14, 19], which is widely available and easy to acquire.
-However, real-world scenarios often require the generation of multi-scene videos from multi-scene descriptions (e.g., Scene1: ‘A koala is napping on a tree.’ Scene2: ‘The koala eats leaves on the tree.’).
-In such cases, the generated video should accurately depict the events in their temporal order (e.g., Scene2 follows Scene1) while maintaining visual consistency, meaning that backgrounds and entities should remain consistent across scenes.
-While high-performance text-to-video models such as Sora [4] might be able to generate multi-scene videos, we point out that they are closed-source models trained with massive compute resources and lack sufficient details on the model design, training protocol, and datasets.
-In this work, we present a complementary approach and tackle the challenge of effectively leveraging the capabilities of base T2V models for multi-scene video generation.
-The multi-scene text-to-video generation differs from long video synthesis where the goal is to either interpolate (few frames to many frames) [8] or create continuing patterns of the single event in the generated video [11].
-Prior works [20, 9] use a transformers [21, 22] to generate video frames for a given scene autoregressively.
-However, it is hard for their model to generate multiple scenes reliably as the context length increases with history of text descriptions and visual tokens [23] of the previous generated videos (e.g., generating Scene 4 conditioned on the Scene1, 2, 3 videos and descriptions).
-Other works [24] utilize a latent diffusion model [25] to generate video frames autoregressively by conditioning on the entire history of generated videos and scene descriptions.
-However, the approach is (a) slow due to repeated sampling, (b) generates only one frame per scene description, and (c) shown to work with only limited cartoon characters [26, 27] instead of wide range of visual concepts in the real-world.
-In this work, our goal is to generate multi-scene videos in the end-to-end manner, using a diffusion text-to-video generative model that is capable of producing content for a wide range of visual entities and actions.
-Figure 1: Multi-scene video generation methods.
-(a) Generating a video by merging scene 1 and scene 2 descriptions.
-(b) The resulting video is composed from the video generated by the description of scene 1 and the video generated by the description of scene 2.
-(c) In our method (TALC) the generated video is conditioned on the description of scene 1 for the first half of the video frames and on the description of scene 2 for the later video frames.
-Figure 2: Examples of multi-scene video generation baselines.
-(a) Generating video on the merged descriptions, leads to a poor text-video alignment.
-(b) Generating videos for the individual text descriptions and concatenate them temporally, leads to a lack of background consistency.
-(c) Our approach (TALC) enhances the scene-level text-video alignment and maintains background consistency.
-As shown in Figure 1(a), the naive approach to generating a multi-scene video for the scene descriptions $(T^{\prime}_{1},T^{\prime}_{2})$ would condition the T2V generative model on the merged descriptions.
-In this setup, the diffusion model processes the entire scene description together, and lacks any information regarding the expected temporal order of events in the generated videos.
-As a result, we find that this approach leads to poor text-video alignment.
-As shown in Figure 1(b), an alternative approach generates videos for the individual text descriptions independently and concatenates them in the raw input space along the temporal dimension.
-While this approach achieves good alignment between the scene description and the scene-specific video segment, the resulting video lacks visual consistency in terms of entity and background appearances.
-Prior work [28, 29] generates multi-scene videos by utilizing knowledge of the entity, background, and their movements from large language models [30].
+In this regard, text-to-video (T2V) models have seen rapid advancements by pretraining on internet-scale datasets of images, videos, and texts [5, 6, 7, 8, 9].
+Specifically, latest T2V models such as ModelScope [10], Lumiere [11], VideoCrafter [12, 13], and OpenSora [14], train a diffusion model [15] with single-scene video-text datasets [16, 17] that are widely available and easy to acquire.
+As a result, these models can only generate single scene videos that depicts a single action or event [^1] While high-performance T2V models such as Sora [4] might be able to generate multi-scene videos, we point out that they are closed-source models trained with massive compute resources and lack sufficient details on the model design, training protocol, and datasets..
+However, real-world scenarios (e.g., movie and ads creation) often require the generation of multi-scene video for text descriptions (e.g., Scene1: ‘A koala is napping on a tree.’ Scene2: ‘The koala eats leaves on the tree.’).
+In such cases, the generated video should accurately depict the events mentioned in the multi-scene script while maintaining visual consistency (e.g., entity and background consistency).
+Multi-scene T2V generation differs from long video synthesis where the goal is to either interpolate (few frames to many frames) [7] or create continuing patterns of the single event in the generated video [8].
+Prior works [18, 19] use a transformer [20, 21] to generate video frames for a given scene autoregressively.
+However, these models straggle to generate multiple scenes reliably as the context length increases with the history of text descriptions and visual tokens [22] of the previous generated videos, e.g., generating Scene 4 conditioned on Scene1, 2, 3 videos and descriptions.
+Some works [23] utilize a latent diffusion model [24] to generate video frames autoregressively by conditioning on the entire history of generated videos and scene descriptions.
+However, this approach is slow due to repeated sampling, generates only one frame per scene description, and is shown to work with only limited cartoon characters [25, 26] and not with a wide range of visual concepts in the real-world.
+Other works [27, 28] generate multi-scene videos by utilizing knowledge of the entity, background, and their movements from large language models [29].
 However, these videos are generated independently for each scene before being merged.
 Moreover, these methods do not offer a way to learn from real-world multi-scene video-text data.
-To remedy these challenges, we propose TALC (Time-ALigned Captions), a simple and effective framework to generate consistent and faithful multi-scene videos.
-As shown in Figure 1(c), our approach conditions the T2V generative model with the knowledge of the temporal alignment between the parts of the multi-scene video and multi-scene descriptions.
-Specifically, TALC conditions the visual representations of earlier video frames on the embeddings of the earlier scene description, and likewise, it conditions the representations of later video frames on the embeddings of the later scene description in the temporal dimension.
-Additionally, the temporal modules in the T2V diffusion architecture allows information sharing between video frames (the first half and the second half) to maintain visual consistency.
+To address these challenges, we propose TALC (Time-ALigned Captions), a simple and effective framework to generate consistent multi-scene videos by effectively leveraging the capabilities of pretrained T2V diffusion models.
+As shown in Figure 1(c), TALC conditions the visual representations of earlier video frames on the embeddings of the earlier scene description, and likewise, it conditions the representations of later video frames on the embeddings of the later scene description in the temporal dimension.
+The temporal modules in the T2V diffusion architecture allows information sharing between video frames (the first half and the second half) to maintain visual consistency.
 Thus, TALC enhances the scene-level text-video alignment while providing all the scene descriptions to the diffusion model at once.
 Further, our TALC framework can enhance the multi-scene text-to-video generation capabilities with real-world multi-scene data (§3.2).
-In our experiments, we assess the visual consistency (background and entity consistency) and multi-scene script adherence of the generated videos from Modelscope [10] and Lumiere [6].
-Through our automatic and human evaluation, we find that merging scene descriptions leads to high visual consistency but poor text adherence.
-On the other hand, we observe that merging videos independently achieves the highest text adherence while the visual consistency is compromised.
-Interestingly, switching to TALC strikes an effective balance between visual consistency and text adherence, outperforming the baseline methods by $11.1$ points on the overall score.
-This score represents the average of visual consistency and text adherence scores, as determined by human evaluation.
+We provide a qualitative example for comparing different multi-scene video generation methods in Appendix §C.
+In our experiments, we assessed the visual consistency (background and entity consistency) and multi-scene script adherence of the generated videos from Modelscope [10] and Lumiere [11].
+In addition, to our TALC method, we evaluated concatenation approaches, either merging descriptions of different scenes or generating a video per scene independently and concating the output videos (Figure 1(a,b)).
+Through our automatic and human evaluation, we found that merging scene descriptions leads to high visual consistency but poor text adherence.
+On the other hand, merging videos independently achieves high text adherence but visual consistency is compromised.
+Our TALC model strikes an effective balance between visual consistency and text adherence, outperforming the baseline methods by $11.1$ points on the overall score, averaging visual consistency and text adherence scores, as determined by human evaluation.
 Furthermore, we construct a multi-scene text-video dataset from real-world videos and fine-tune the T2V generative model using TALC.
@@
 2 Preliminaries
-In this work, we focus on generating multi-scene videos from scene descriptions using a diffusion-based Text-to-Video (T2V) generative model.
+In this work, we focus on generating multi-scene videos from text descriptions using a diffusion-based Text-to-Video (T2V) generative model.
 The initial step is to equip the generative model with the knowledge of a wide range of visual concepts and actions.
 This is achieved during the pretraining stage (§2.1).
-Subsequently, we aim to utilize the base model for multi-scene text-to-video generation task, which we formalize in (§2.3).
-In §3, we propose our TALC framework and discuss collection of real-world multi-scene text-video data for finetuning the base T2V model.
+Subsequently, we aim to utilize the pretrained model for multi-scene text-to-video generation task, which we formalize in (§2.3).
 2.1 Diffusion Models for Text-to-Video Generation
 Diffusion models [15, 30] $p_{\theta}(x)$ are a class of generative models that learn data distribution $p_{data}(x)$.
-Due to their flexible design, we can train their class-conditional versions to learn class-conditional data distributions $p_{data}(x|y)$ where $y$ is the conditioning variable, that can take various forms such as labels from a dataset or text description accompanying in a video [32].
+Their class-conditional versions can be trained to learn conditional distributions $p_{data}(x|y)$ where $y$ is the conditioning variable, which could take various forms such as labels from a dataset or text description accompanying a video [31].
 We assume a dataset $\mathcal{S}\subset\mathcal{V}\times\mathcal{T}$ consisting of pairs of $(V_{j},T_{j})$ where $V_{j}\in\mathbb{R}^{L\times 3\times H\times W}$ is a raw video consisting of $3$ RGB channels, $L$ frames, $H$ height, $W$ width, and $T_{j}$ is a text caption.
-We use $\mathcal{V}$ and $\mathcal{T}$ to denote the domain of videos and text, respectively.
-The aim of T2V generative modeling is to learn the conditional distribution of the videos conditioned on the text $p_{\mathcal{S}}(V_{j}|T_{j})$.
-In this work, we consider diffusion-based generative models that learn the data distribution via iterative denoising of the input video $z_{j}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$.
-Here, $z_{j}$ can either represent the input video in the raw pixel space $V_{j}$ [6] or it can represent the latent representation of the video $z_{j}=\mathcal{E}(V_{j})$ for the latent diffusion models [25] where $\mathcal{E}$ is an encoder network such as VAE [33].
-Given $z_{j}$, diffused variable $z_{\tau,j}=\alpha_{\tau}z_{j}+\beta_{\tau}\epsilon$ are constructed where $\epsilon\sim\mathcal{N}(0,I)$ where $\alpha_{\tau}$ and $\beta_{\tau}$ are sampled from the noise scheduler $p_{\tau}$ [34] which define the noise levels the model is trained on.
-Finally, we train a denoiser network $f_{\theta}$ [35, 36] that inputs the diffused variable $z_{\tau}$ and embeddings of the text caption to predict the target vector $y$ where $y$ can be the original noise $\epsilon$, which minimizes the denoising score matching objective [13]:
+The T2V model is trained to learn the conditional distribution of the videos conditioned on the text $p_{\mathcal{S}}(V_{j}|T_{j})$.
+In this work, we consider diffusion-based generative models that learn this distribution via iterative denoising of the video $z_{j}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$.
+Here, $z_{j}$ can either represent the video in the raw pixel space $V_{j}$ [11] or it can represent the latent representation of the video $z_{j}=\mathcal{E}(V_{j})$ for the latent diffusion models [24] where $\mathcal{E}$ is an encoder network such as VAE [32].
+Given $z_{j}$, diffused variable $z_{\tau,j}=\alpha_{\tau}z_{j}+\beta_{\tau}\epsilon$ are constructed where $\epsilon\sim\mathcal{N}(0,I)$ and $\alpha_{\tau}$ and $\beta_{\tau}$ are sampled from the noise scheduler $p_{\tau}$ [33], which defines the noise levels the model is trained on.
+Finally, a denoiser network $f_{\theta}$ [34, 35], which takes the diffused variable $z_{\tau}$ and embeddings of the text caption as input, is trained to predict the target vector $y$ where $y$ can be the original noise $\epsilon$, which minimizes the denoising score matching objective [15]:
 $\mathbb{E}_{(V_{j},T_{j})\in S,\tau\sim p_{\tau},\epsilon\sim N(0,I)}\left[||\epsilon-f_{\theta}(\tau,z_{\tau,j},h_{j})||_{2}^{2}\right]$ (1)
-where $h_{j}=\mathcal{H}(T_{j})\in\mathbb{R}^{d}$ is the embedding of the text caption $T_{j}$ where $\mathcal{H}$ is the text embedding model [37] and $d$ is the dimension size.
+where $h_{j}=\mathcal{H}(T_{j})\in\mathbb{R}^{d}$ is the embedding of the text caption $T_{j}$, $\mathcal{H}$ is the text embedding model [36] and $d$ is the dimension size.
 2.2 Text Conditioning Mechanism
-To ensure the effective textual controllability of video generation, the structure of the denoiser networks is equipped with a cross-attention mechanism [10, 8].
+For effective textual controllability of video generation, the denoiser networks is equipped with a cross-attention mechanism [10, 7].
 Specifically, it conditions the visual content $z_{\tau}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ on the text.
@@
 These repeated text embeddings $r_{j}$ are inputs to the spatial attention block as the key and value in the multi-head attention block.
-The cross-attention enables the intermediate visual features to capture the semantic information that facilitates an alignment between the language and vision embeddings.
+The cross-attention enables the intermediate visual features to capture the semantic information that facilitates an alignment between the text and vision embeddings.
 Formally,
 $z^{\prime}_{\tau,j}=CA_{f_{\theta}}(Q=z_{\tau,j};K=r_{j};V=r_{j})$ (2)
-where $CA_{f_{\theta}}$ is the cross attention mechanism with $Q,K,V$ as the query, key, and value, respectively, in the spatial blocks of the denoiser network.
-Additionally, $z^{\prime}_{\tau,j}$ is the intermediate representation that is informed with the visual and textual content of the data.
-In addition to the spatial blocks, the denoiser network also consists temporal blocks that aggregate features across video frames which are useful for maintaining visual consistency in the generated video.
+where $CA_{f_{\theta}}$ is the cross attention mechanism with $Q,K,V$ as the query, key, and value, respectively, in the spatial blocks of the denoiser network, $z^{\prime}_{\tau,j}$ is the intermediate representation that is informed with the visual and textual content of the data.
+In addition to the spatial blocks, the denoiser network also consists of temporal blocks that aggregate features across video frames, which are useful for maintaining visual consistency in the generated video.
 2.3 Multi-Scene Text-to-Video Generation
-In many real-world scenarios, such as movies, stories, and instructional videos [38], a video may depict multiple transitions with the same or changing entities, as well as multiple actions or events.
-In addition, the different video segments often share contextual information such as the background or location.
-These videos are considered multi-scene videos.
-In this work, we aim to generate multi-scene video $X=\{x_{1},x_{2},\ldots,x_{n}\}$ from multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ where $n$ are the number of sentences and each sentence $y_{j}$ is a scene description for scene $j$.
-Additionally, the index $j$ also defines the temporal order of events in the multi-scene script i.e., we want the events described in the scene $j$ to be depicted earlier than the events described in the scene $k$ where $k>j$.
-Further, we want the parts of the entire generated video $X$, given by $x_{j}$, to have high video-text semantic alignment with the corresponding scene description $y_{j}$, also referred to as text adherence.
-For instance, consider a two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’}.
-Here, we need the T2V generative model to synthesize the appearance of the red panda (an entity) that remains consistent throughout the generated video, also referred to as entity consistency.
-In addition, we will expect that the context of the multi-scene video of a forest (a background) to remain consistent, also referred to as background consistency.
-Figure 3: The architecture of Time-Aligned Captions (TALC).
-During the generation process of the video, the initial half of the video frames are conditioned on the embeddings of the description of scene 1 ($r_{y_{1}}$), while the subsequent video frames are conditioned on the embeddings of the description of scene 2 ($r_{y_{2}}$).
-3 Method
-3.1 TALC: Time-Aligned Captions for Multi-Scene T2V Generation
-Most of the existing T2V generative models [10, 16, 6] are trained with large-scale short video-text datasets (10 seconds - 30 seconds) such as WebVid-10M [14].
-Here, each instance of the dataset consists of a video and a human-written video description.
-These videos either lack the depiction of multiple events, or the video descriptions do not cover the broad set of events in the video, instead focusing on the major event shown.
-As a result, the pretrained T2V generative models only synthesize single video scenes depicting individual events.
-We introduce TALC, a novel and effective framework to generate multi-scene videos from diffusion T2V generative models based on the scene descriptions.
-Our approach focuses on the role of text conditioning mechanism that is widely used in the modern T2V generative models (§2.2).
-Specifically, we take inspiration from the fact that the parts of the generated video $x_{j}$ should depict the events described in the scene description $y_{j}$.
-To achieve this, we ensure that the representations for the part of the generated video aggregates language features from the scene description $y_{j}$.
-Consider that we want to generate a multi-scene video $X\in\mathbb{R}^{L\times 3\times H\times W}$ from the scene descriptions $y_{j}\in Y$, using a T2V generative model $f_{\theta}$.
-Furthermore, we assume that individual video segments $x_{j}$ are allocated $L/n$ frames within the entire video $X$.
-Let $z_{X}=[z_{x_{1}};z_{x_{2}};\ldots;z_{x_{n}}]\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ represent the representation for the entire video $X$, and $z_{x_{j}}\in\mathbb{R}^{(L/n)\times C\times H^{\prime}\times W^{\prime}}$ for the $j^{th}$ part of the video that are concatenated in the temporal dimension.
-In addition, consider $r_{Y}=\{r_{y_{1}},\ldots,r_{y_{n}}\}$ be the set of text embeddings for the multi-scene description $Y$ and $y_{j}$ be an individual scene description.
-In the TALC framework, the Eq.
+We aim to generate multi-scene video $X=\{x_{1},x_{2},\ldots,x_{n}\}$ from multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ where $n$ are the number of scenes and each sentence $y_{j}$ is a description for scene $j$.
+Index $j$ also defines the temporal order of events in the multi-scene script, i.e., if $k>j$, we want the events described in scene $j$ to be depicted earlier than the events described in scene $k$.
+Further, we want that each video part $x_{j}$, within the entire generated video $X$, to have high video-text semantic alignment with its corresponding scene description $y_{j}$.
+Figure 2: Time-Aligned Captions (TALC) Framework.
+During the generation process of the video, the initial half of the video frames are conditioned on the embeddings of the description of scene 1, while the subsequent video frames are conditioned on the embeddings of the description of scene 2.
+3 TALC: Time-Aligned Captions for Multi-Scene T2V Generation
+In this section, we present our novel method TALC, describing its architecture and a novel approach for constructing synthetic multi-scene video-text data to train TALC.
+3.1 TALC Architecture
+Our starting point is that each specific scene $x_{j}$ in a multi-scene video $X\in\mathbb{R}^{L\times 3\times H\times W}$ should reflect the entities and events described in the conditioning scene description $y_{j}$.
+To achieve this, TALC adapts T2V diffusion architecture $f_{\theta}$ to multi-scene generation by ensuring that the representation for video part $x_{j}$ attends and aggregates language features specifically from description $y_{j}$.
+Concretely, TALC allocates $L/n$ frames within the entire video to each $x_{j}\in X$.
+Let $z_{X}=[z_{x_{1}};z_{x_{2}};\ldots;z_{x_{n}}]\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ be the representation for the entire video $X$, and $z_{x_{j}}\in\mathbb{R}^{(L/n)\times C\times H^{\prime}\times W^{\prime}}$ for the different scenes, which are concatenated in the temporal dimension.
+In addition, $r_{Y}=\{r_{y_{1}},\ldots,r_{y_{n}}\}$ be the sequence of text embeddings for the multi-scene description $Y$ and $y_{j}$ be an individual scene description.
+Under TALC, Eq.
 2 is changed to:
@@
 $z^{\prime}_{\tau,X}$ $=[z^{\prime}_{x_{1}};z^{\prime}_{x_{2}};\ldots;z^{\prime}_{x_{n}}]$ (4)
-Here, $\tau$ represents the timestamp in the diffusion modeling setup, which is applied during training as well as inference.
-We illustrate the framework in Figure 3.
-While TALC aims to equip the generative model with the ability to depict all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the denoiser network.
+with, $\tau$ the timestamp in the diffusion modeling setup, which is applied during training as well as inference. We illustrate the framework in Figure 2.
+While TALC aims to equip the generative model with the ability to follow all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the existing denoiser network.
 By design, our approach can be applied to the pretrained T2V model during inference.
-3.2 Baselines
-Here, we describe the baseline methods that could be used to generate videos for the multi-scene descriptions from a given diffusion text-to-video generative model.
-3.2.1 Merging Captions
-In this setup, we create a single caption by merging all the multi-scene descriptions.
-Specifically, the multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ can be written as a single prompt ‘$\mathcal{P}=y_{1}.\text{Then, }y_{2}.\ldots\text{Then, }y_{n}.$’ For instance, the two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’} will change to $\mathcal{P}=$ ‘A red panda climbs on a bamboo forest.
-Then, the red panda sleeps peacefully in the treetop.’ Subsequently, we generate a video from the T2V model $f_{\theta}$ by conditioning it on $\mathcal{P}$.
-While this approach mentions the temporal sequence of the events in a single prompt, the T2V model does not understand the temporal boundaries between the two events.
-Specifically, the Eq.
-2 suggests that the visual features for all the video frames will aggregate information from the entire multi-scene description, at once, without any knowledge about the alignment between the scene description and its expected appearance in the generated video.
-3.2.2 Merging Videos
-In this setup, we generate videos for each scene description individually and merge them in the raw input space.
-Formally, the individual scene description $y_{i}$ conditions the T2V model $f_{\theta}$ to generate the parts of the multi-video $x_{i}$.
-Finally, we stitch the individual videos together to synthesize the entire video $X={x_{1},x_{2},\ldots,x_{n}}$.
-In this process, the parts of the multi-scene video closely adhere to the scene descriptions, leading to high text fidelity.
-However, since the generated videos do not have access to all the multi-scene descriptions (e.g., the video for Scene 2 is not informed about Scene 1), the visual consistency across the entire video is quite poor.
-3.3 Multi-Scene Video-Text Data Generation
+3.2 TALC Training
 Figure 3: Our approach for generating time-aligned video captions.
@@
 These frames are processed collectively through the Gemini model to produce multi-image captions that maintain narrative continuity by contextualizing each scene within the video’s overall sequence.
-While our approach generates better multi-scene videos, the text adherence capabilities of the pretrained T2V generative model are limited.
-This is due to the lack of multi-scene video-text data during its pretraining.
-Unlike single video-text datasets, the multi-scene video-text datasets are not widely available and are hard to curate for model training.
-This is attributed to the fact that high-quality caption generation requires a lot of human labor which is time-consuming and expensive.
-Prior work such as ActivityNet [39] has curated human captions for specific video scenes depicting useful actions in long videos.
-However, the video scenes are either overlapping or have a large temporal gap between them that will be harmful for natural and smooth variations between the generated multi-scene videos.
-Hence, the absence of high-quality captions for continuous video scenes in the dataset makes unsuitable for T2V generative training.
-To this end, we aim to create a real-world multi-scene video-text dataset to allow further training of the pretrained T2V models.
-Specifically, we leverage the capability of the multimodal foundation model, Gemini-Pro-Vision [40], to generate high-quality synthetic data for enhanced video-text training [41].
-Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description from the dataset.
-Subsequently, we utilize PySceneDetect library  [^1] https://github.com/Breakthrough/PySceneDetect to generate continuous video scenes from $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of scene cuts in the video.
-A similar approach was used in a prior work [12] to detect scene changes in the video data.
-Then, we sample the middle video frame $F_{i,j}$ as a representative of the semantic content in the video scene $A_{i,j}$.
-Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for a single video $A_{i}$ and the entire video caption $B_{i}$ to a large multimodal model [40].
-Specifically, the model is prompted to generate high-quality captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$.
-We provide the prompt provided to the multimodal model in Appendix §A.
-In Figure 4 we provide an instance for the multi-scene video-text data generation.
-Datasets.
-To construct a multi-scene video-text dataset, we utilize existing dataset that include natural (real) videos and associated high-quality human-written captions that summarize the entire video.
-Specifically, we choose MSR-VTT [42] and VaTeX [43].
-Most of the videos in MSR-VTT are 10-30 seconds long while VaTeX consists 10 seconds long videos.
-In addition, each video in MSR-VTT and VaTex consists 20 captions and 10 captions, respectively, out of which one is selected at random for multi-scene data generation.
-As described above, a single video is cut into multiple video segments using Pyscene library.
-In our experiments, we retain the first four video segments and discard any additional segments if the library generates more than four.
-Since the accuracy of the multi-scene captioning and the computational demands during finetuning are influenced by the number of scenes, we opt to limit the scene count to four for our experiments.
-However, future work could employ similar methodologies to scale the number of scenes, given more computing power and advanced multi-scene captioning models.
-We provide the data statistics for the final multi-scene data in Appendix §G.
-4 Evaluation
-In this section, we describe the evaluation scheme for videos generated from multi-scene text descriptions.
-First, we describe the evaluation metrics that we aim to assess in this work (§4.1).
-Then, we generate multi-scene descriptions for a diverse set of tasks (§4.2).
-Finally, we present the details for automatic and human evaluation of the generated videos (§4.3).
-4.1 Metrics
-The ability to assess the quality of the generated multi-scene videos is a challenging task itself.
-As humans, we can judge the multi-scene videos across diverse perceptual dimensions [44] that the existing automatic methods often fails to capture [45].
-Following [28], we focus on the visual consistency of the generated video, text adherence capabilities of the T2V models, and video quality of the video.
-Here, we present the metrics with the aspects that they intend to assess in the generated video for multi-scene text description.
-Visual Consistency.
-This metric aims to assess the (entity or background) consistency between the frames of the multi-scene videos.
-Here, the entity consistency aims to test whether the entities in the multi-scene video are consistent across the video frames.
-For instance, the appearance of an animal should not change without a change described in the text description.
-In addition, the background consistency aims to test whether the background of the multi-scene video remains consistent across the video frames.
-For instance, the room should not change without a change description in the text.
-Text Adherence.
-This metric aims to test whether the generated video adheres to the multi-scene text description.
-For instance, the events and actions described in the text script should be presented in the video accurately, and in the correct temporal order.
-In our experiments, we compute the visual consistency and text adherence with the automatic and human evaluators.
-Further, we compute the overall score, which is the average of the visual consistency and text adherence scores.
-In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Table 1.
-4.2 Task Prompts
-Here, we curate a set of task prompts for diverse scenarios, aiming to holistically assess the quality of the generated videos.
-Single character in multiple visual contexts (S1).
-In this scenario, we instruct an LLM, GPT-4, to create a coherent script consisting of four scenes.
+We note that the text adherence capabilities of the pretrained T2V generative model are limited.
+This is due to the lack of multi-scene video-text data during pretraining, since these datasets are not widely available and are hard to curate for model training.
+Prior work, such as ActivityNet [37], curated human captions for specific video scenes depicting useful actions in long videos.
+However, these scenes are either overlapping or have a large temporal gap between them, which are less suitable for natural and smooth transitions required in multi-scene videos.
+To this end, we propose a method for creating synthetic real-world multi-scene video-text examples, with which we further train our T2V model.
+Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description.
+Then, similar to [9], we apply PySceneDetect [^2] https://github.com/Breakthrough/PySceneDetect to segment $A_{i}$ into continuous video scenes $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of detected scenes.
+Next, we sample the middle video frame $F_{i,j}$ in each scene $A_{i,j}$ as its semantic representative.
+Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for $A_{i}$ and the entire video caption $B_{i}$ to Gemini-Pro-Vision-1.0 [38].
+Gemini is prompted to generate captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$.
+In Figure 3, we provide an instance for the multi-scene video-text data generation.
+We provide the prompt provided to Gemini and the source datasets in Appendix §D and §E, respectively.
+4 Experimental Setting
+In this section, we describe the setup of our experiments for generated multi-scene videos, including eval-set creation, metrics, automatic and human evaluation, and the tested models.
+4.1 Eval-Set Task Prompts
+To create an eval-set for testing models that generate multi-scene videos, we curate a set of task prompts for diverse scenarios (The full task prompts are in Appendix §F.).
+We generate $100$ prompts for each scenario.
+Single character in multiple visual contexts (S1): We instruct GPT-4 [39] to create a coherent script consisting of four scenes.
 Each scene features a specific animal character performing diverse activities in every scene.
-This task assesses the capability of the T2V model to generate consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script.
-For instance, a generated script could be ‘Scene 1: A red panda is climbing a tree.
+This task assesses the consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script.
+An example script is “Scene 1: A red panda is climbing a tree.
 Scene 2: The red panda eats the leaves on the tree.
-Scene 3: The red panda lies down on the branch of the tree.
-Scene 4: The red panda sleeps on the branch’.
-In total, we generate $100$ prompts in this scenario.
-Different characters in a specific visual context (S2).
-In this scenario, we instruct a language model, GPT-4, to create a coherent script consisting of four scenes.
-Each scene features different animal characters engaging in the same activity in every scene [20].
-This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the appearance of the different characters in the multi-scene text script.
-For instance, a generated script could be ‘Scene 1: A cat leaps onto countertop.
+Scene 3: … Scene 4: …”.
+Different characters in a specific visual context (S2): We instruct GPT-4 to create a coherent script consisting of four scenes.
+Each scene features the same activity but with different animal characters [18].
+This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the animal character changes.
+An example script is “Scene 1: A cat leaps onto a countertop.
 Scene 2: A dog leaps onto the same countertop.
-Scene 3: A rabbit leaps onto the same countertop.
-Scene 4: A raccoon leaps onto the same countertop’.
-In total, we generate $100$ prompts in this scenario.
-Multi-scene captions from real videos (S3).
-Here, we aim to assess the ability of the model to generate multi-scene videos for open-ended prompts that are derived from real-world videos.
-This task also assesses the ability of the T2V model to generate consistent appearances of the entity and its background while adhering to multi-scene descriptions.
-Specifically, we use our multi-scene video-text data generation pipeline (§3.3) to create such prompts for the real videos from the test splits of the video-text datasets.
-For example, a multi-scene text script could be ‘Scene 1: A beauty vlogger introduces her skincare routine.
-Scene 2: She applies a serum to her face, smoothing it in’.
-We present a sample of the various task prompts in the Appendix §B.
-In total, we generate $100$ prompts in this scenario.
-4.3 Evaluator
-In this work, we devise an automatic evaluation framework and perform human evaluation to assess the quality of the multi-scene generated videos.
-Automatic Evaluation.
-Here, we utilize the capability of a large multimodal model, GPT-4-Vision [46], to reason over multiple image sequences.
-First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for two-scene video).
-Then, we prompt the multimodal model with the temporal sequence of video frames from different scenes and the multi-scene text description.
-Specifically, we instruct the multimodal model to decide the quality of the generated video across various metrics including entity consistency, background consistency, and text adherence.
-For each metric, the multimodal model assigns one of three possible response $\{yes=1,partial=0.5,no=0\}$.
-For instance, $yes$ for the entity consistency metric implies that the video frames sampled from the generated video have consistent appearance of the entity described in the multi-scene script.
-In this work, we do not utilize any existing video-text alignment models [47, 41] for evaluating text adherence as they are trained on single-scene video-text datasets.
-We present the automatic evaluation prompt in Appendix §C.
-Human Evaluation.
-We also conduct a human evaluation to assess the multi-scene generated videos along the dimensions of visual consistency, text adherence, and visual quality.
-Specifically, we ask the annotators from Amazon Mechanical Turk (AMT) to choose one of three options for each metric $\{yes,partial,no\}$, similar to the automatic evaluation.
-In addition, we choose the annotators that pass a preliminary qualification exam.
-We present the screenshot of the UI in Appendix §D.
-4.4 Evaluation Setup
-Since merging captions (§3.2) and TALC3.1) methods input the entire multi-scene text description at once, the quality of the video generated by these methods is influenced by the number of scenes described in the text script.
-Hence, we calculate the performance of the baselines and TALC by averaging the scores assigned to videos generated for two, three, and four scenes.
-Additionally, we report on visual consistency by averaging the performance across the entity and background consistency metrics.
-Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), and the background consistency and text adherence scores are computed for all the task prompts.
-We also evaluate the impact of TALC-based finetuning on the single scene generation in Appendix §I.
+Scene 3: A rabbit … Scene 4: A raccoon …”.
+Multi-scene captions from real videos (S3): We use our multi-scene video/text data generation pipeline (§3.2) to create similar prompts for real videos for the test splits of the video-text datasets.
+This task assesses the ability of a model to generate multi-scene videos with consistent appearances of entities and background for open-ended prompts, which are derived from real videos.
+An example script is “Scene 1: A vlogger introduces her skincare routine.
+Scene 2: She applies a serum to her face.
+Scene 3: …”.
+4.2 Evaluation Procedure
+Following [27], we assess the capability of the tested T2V models to: (a) follow the text instructions in the script (Text Adherence); (b) keep consistent appearance of entities and background across scenes (Visual Consistency), unless the script specifically requests a change in any of them.
+We compute the visual consistency and text adherence with automatic and human evaluators.
+We also report an overall score, which is the average of the visual consistency and text adherence scores.
+In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Appendix M):
+Automatic Evaluation: We utilize GPT-4-Vision (GPT4V) [40] to reason over multiple images.
+First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for a two-scene video).
+Then, we prompt GPT4V with the temporal sequence of the sampled frames and the multi-scene text description and instruct GPT4V to decide on the quality of the generated video across several attributes, including entity consistency, background consistency, and text adherence.
+For each attribute, GPT4V assigns one of three possible response $\{\textit{yes}=1,\textit{partial}=0.5,\textit{no}=0\}$.
+The automatic evaluation prompt is given in Appendix §G.
+[…]
+Since merging captions (§4.3.1) and TALC3.1) methods input the entire multi-scene text description at once, the quality of the video generated by these methods is influenced by the number of scenes described in the text script.
+Hence, we calculate the performance of the tested models by averaging the scores assigned to videos generated for two, three, and four scenes, over 900900900900333100 (examples per scenario) * 3 (scenarios) * 4 (2-scene/3-scene/4-scene) = 900.
+multi-scene prompts.
+We report visual consistency by averaging the performance across the entity and background consistency metrics.
+Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), while the background consistency and text adherence are computed for all the task prompts.
+4.3 Tested Models
+We experiment with ModelScope [10] and Lumiere [11] T2V models for multi-scene video generation with TALC.
+We mainly experiment with ModelScope due to its ease of reproduction and adoption in prior work [27].
+We also experiment we Lumiere to showcase TALC’s generalization.
+By design, TALC can be applied to a base T2V model during inference, without any multi-scene finetuning.
+We compare the multi-scene performance of the two base models without finetuning under three settings: TALC, merging captions, and merging videos (see baseline details in Section 4.3.1).
+In addition, we finetune ModelScope with the TALC framework using the multi-scene video-text data (§3.2).
+As a baseline, we also finetune ModelScope without the TALC framework by naively merging the scene-specific captions in the raw text space.
+In both cases, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$.
+We provide more finetuning details in Appendix §L.
+During inference, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere. More inference details are found in Appendix §J.
+4.3.1 Baselines
+For Merging Captions, we create a single caption by merging all the multi-scene descriptions.
+Specifically, the multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ can be written as a single prompt ‘$\mathcal{P}=y_{1}.\text{Then, }y_{2}.\ldots\text{Then, }y_{n}.$’.
+For instance, the two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’} would change to $\mathcal{P}=$A red panda climbs on a bamboo forest.
+Then, the red panda sleeps peacefully in the treetop.” Subsequently, we generate a video from the T2V model $f_{\theta}$ by conditioning it on $\mathcal{P}$.
+For Merging Videos, we generate a video $x_{i}$ for each scene description $y_{i}$ individually and stitch the output videos together afterwards to synthesize the entire video $X={x_{1},x_{2},\ldots,x_{n}}$.
 5 Experiments
-5.1 Text-to-Video Generative Models
-In this work, we utilize ModelScope [10] and Lumiere [6] T2V models for multi-scene video generation.
-Here, ModelScope is an open-source T2V model with 1.7 billion parameters including the video encoder, text encoder, and denoising U-net network.
-Specifically, it is trained to generate 16 video frames on the mix of WebVid [14] video-text dataset and LAION [48] image-text dataset.
-We perform most of our experiments on ModelScope due to its easy-of-access and adoption in prior works [28].
-In addition, we also include Lumiere-T2V, a model that leverages space-time U-Net denoising networks to generate high-quality videos.
-In this work, we include early experiments with Lumiere to showcase the flexibility of the TALC approach for multi-scene video generation.
-Base model with TALC.
-As described in §3.1, our approach modifies the traditional text-conditioning mechanism to be aware of the alignment between text descriptions and individual video scenes.
-By design, the TALC framework can be applied to the base T2V model during inference, without any multi-scene finetuning.
-Thus, we compare the performance of the multi-scene videos generated from ModelScope and Lumiere T2V base models under three settings: merging captions, merging videos, and TALC.
-In this setting, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere.
-We provide more details on the inference in Appendix §F.
-Finetuning with TALC.
-Since the base model is pretrained with single-scene data, we aim to show the usefulness of TALC framework when we have access to the multi-scene video-text data.
-To this end, we finetune ModelScope on the multi-scene video-text data (§3.3) with TALC framework.
-As a pertinent baseline, we also finetune the ModelScope without TALC framework by naively merging the scene-specific captions in the raw text space.
-In this setting, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$.
-We provide further details on the finetuning setup in Appendix §H.
-The inference settings are identical to the prior method of generating videos from the base model without finetuning.
 (a) Performance on ModelScope T2V model.
@@
 We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
-In this section, we present the results for the baselines and TALC framework averaged over a diverse task prompts and multiple scenes using automatic evaluation (§5.2) and human evaluation (§5.3).
-Finally, we provide qualitative examples for the multi-scene generated videos to showcase the usefulness of our approach (§5.4).
-5.2 Automatic Evaluation
+5.1 Automatic Evaluation Results
 We compare the performance of the baselines (e.g., merging captions and merging videos) with the TALC framework for ModelScope and Lumiere using the automatic evaluation in Figure 4.
 TALC outperforms the baselines without any finetuning.
-In Figure 5(a), we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($68.6$ points), outperforms the overall score achieved by the videos generated using merging captions ($61.7$ points) and merging videos ($67.5$ points) with the base ModelScope.
+In Figure 4(a), we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($69$ points), outperforms the overall score achieved by the videos generated using merging captions ($62$ points) and merging videos ($68$ points) with the base ModelScope.
 Specifically, we observe that the visual consistency of the generated video is high for merging captions ($91$ points) and TALC ($89.9$ points) while it is low for merging videos ($65$ points).
 This indicates that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames.
+Since merging captions mentions multiple events in a single prompt, the T2V model struggles to understand the temporal boundaries between these events.
+As indicated by Eq.
+2, the visual features for all video frames aggregate information from the entire multi-scene description simultaneously, without aligning individual scene descriptions with their corresponding appearances in the generated video.
+In Figure 4(a), this results in poor text adherence with merging captions ($32$ points).
 In addition, we observe that the text adherence using TALC outperforms merging captions by $14.8$ points, while the text adherence is the highest with a score of $70$ points using merging videos.
@@
 Further, we find that TALC outperforms merging captions by $5$ points on text adherence, while merging videos achieves the highest text adherence $65$ points.
-This highlights that the model more easily generates multi-scene videos that adhere to individual text scripts, whereas adherence to the text diminishes when the model is given descriptions of multiple scenes all at once.
+This highlights that the TALC framework enhances the adherence to the scene descriptions.
 Finetuning with TALC achieves the best performance.
@@
 To perform fine-grained analysis of the performance, we assess the visual consistency and text adherence scores for the baselines and TALC framework across diverse task prompts and number of scenes on ModelScope.
-We present their results in Appendix §E.
+We present their results in Appendix §I.
 In our analysis, we find that finetuning with TALC achieves the highest overall score over the baselines across all the scenarios.
@@
 Overall, we show that the TALC strikes a good balance between visual consistency and text adherence to generate high-quality multi-scene videos.
-5.3 Human Evaluation
+5.2 Human Evaluation Results
 Figure 5: Human evaluation results for ModelScope model.
@@
 We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
-Table 1: Human evaluation results on the visual quality of the generated videos from ModelScope.
-We observe that the visual quality of the generated videos are close to each other for the base model.
-However, finetuning the model with merging captions reduces the video quality by a large margin while TALC-finetuned model retains the video quality.
-Method Quality
-Merging Captions (Base) 80.5
-Merging Videos (Base) 86.5
-TALC (Base) 84.5
-Merging Captions (F.T.) 63.4
-TALC (F.T.) 83.3
 TALC achieves the best performance in human evaluation.
@@
 Our results highlight at the benefit of including the inductive bias of temporal alignment between the video scenes and their scene descriptions for multi-scene video generation.
-Visual quality of the generated videos.
-We compare the visual quality of the generated videos using human evaluation in Table 1.
-We find that the visual quality of videos generated from the base model ranges from $80.5-86.5$ using the baselines and TALC framework.
-However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points.
-This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video.
-Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining.
-While our work is centered around multi-scene evaluation, we also perform single-scene evaluation in Appendix §I.
+We perform human evaluation on the visual quality of the generated multi-scene videos in §M.
+While our work is centered around multi-scene evaluation, we also perform single-scene evaluation in Appendix §N.
 5.3 Qualitative Analysis
@@
 Our analysis reveals that both ModelScope and Lumiere are capable of producing multi-scene videos that exhibit high text adherence and visual consistency.
-Considering the case of the same animal engaging in multiple actions (referred to as "one character multiple contexts").
+Considering the case of the same animal engaging in multiple actions (referred to as ‘one character multiple contexts’).
 The videos generated by ModelScope successfully maintained the same animal while varying the background and action between the scenes.
@@
 We believe that this distinction is attributed to ModelScope’s fine-tuning with TALC.
-Considering different animals within a particular visual setting (referred to as "multiple-characters same context"), both ModelScope and Lumiere demonstrated impressive abilities in preserving the consistency of the background across the videos and adhering closely to the provided text.
-During our analysis, we noticed that the multi-scene captions derived from real videos (referred to as "open-ended captions") exhibited a substantial number of changes between the various scenes.
+Considering different animals within a particular visual setting (referred to as ‘multiple-characters same context’), both ModelScope and Lumiere demonstrated impressive abilities in preserving the consistency of the background across the videos and adhering closely to the provided text.
+During our analysis, we noticed that the multi-scene captions derived from real videos (referred to as ‘open-ended captions’) exhibited a substantial number of changes between the various scenes.
 In this scenario, Lumiere, when employed without fine-tuning, displayed challenges in adhering to the text, while ModelScope achieved a higher degree of text adherence but was also prone to visual artifacts.
 6 Related Work
+Text-to-Video Generative Modeling.
+The field of text-to-video (T2V) synthesis has significantly evolved from its inception with models like VGAN [2] and MoCoGAN [41], leveraging the foundational technologies of GANs [42] and VAEs [43] to produce concise, single-scene videos.
+The advent of diffusion models, exemplified by Imagen Video [44], marked a notable advancement in T2V synthesis.
+Despite these strides, the challenge of creating multi-scene videos that reflect the complexity of the physical world [1, 2, 3] remains.
+Our work, TALC, extends the capabilities of T2V models to multi-scene storytelling, filling a crucial gap in the synthesis landscape.
+Multi-Scene Video Generation.
+The pursuit of multi-scene T2V synthesis has been furthered by recent innovations like Phenaki [18] and Stable Video Diffusion [9], which have explored new frontiers in video generation from textual prompts and the scaling of latent diffusion models, respectively.
+Despite these advancements, the task of generating videos that convey coherent narratives across multiple scenes remains formidable, highlighted by recent works such as VideoPoet [19].
+TALC tackles this task and offers a framework produces videos spanning multiple scenes.
+We also introduce nuanced evaluation approach.
+This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis.
+We provide a detailed related work in §B.
+7 Conclusion
+We introduced TALC, a simple and effective method for improving the text-to-video (T2V) models for multi-scene generation.
+Specifically, it incorporates the knowledge of the natural alignment between the video segments and the scene-specific descriptions.
+Further, we show that TALC-finetuned T2V model achieve high visual consistency and text adherence while the baselines suffer from one or both of the metrics.
+Given its design, our framework can be easily adapted into any diffusion-based T2V model.
+An important future direction will be to scale the amount of multi-scene video-text data and deploy TALC framework during pretraining of the T2V models.
+8 Acknowledgement
+We would like to thank Ashima Suvarna for providing feedback on the draft. Hritik Bansal is supported in part by AFOSR MURI grant FA9550-22-1-0380.
+Appendix A Limitations
+While there are numerous text-to-video generative models, our work focuses on diffusion-based T2V models due to their higher visual quality and better latency in comparison to autoregressive models [19].
+In our work, we argue that the text conditioning mechanism is unified across diffusion T2V models, and thus TALC framework should generalize to various model architectures.
+Specifically, we perform our experiments with ModelScope, a latent diffusion model, and Lumiere, a diffusion model that operates in the raw video space.
+In the future, it will helpful to adopt TALC for multi-scene video generation in novel training and model design paradigms.
+In this work, we evaluate the multi-scene video generation capabilities uptill four scenes.
+This is at par with the number of scenes in prior multi-scene video generation works [18, 23].
+In our experiments, we further show that the performance of the models degrade as we generate more scenes, and hence, we limited ourselves to four scenes.
+A relevant future direction would be to expand the quality of multi-scene video-text data beyond four scenes and finetune T2V models using TALC framework.
+In addition, we perform human annotations using Amazon Mechanical Turkers (AMT) where most of the workers primarily belong to the US and Canada (hourly rate - \$$\$18$ per hour).
+Hence, the human annotations in this work do not represent the diverse demographics around the globe.
+As a result, our human annotations reflect the perceptual biases of the annotators from Western cultures.
+In the future, it will be pertinent to assess the impact of diverse groups on our human evaluations.
+Finally, we acknowledge that the text-to-video generative models can perpetrate societal biases in their generated content [45, 46].
+It is critical that future work quantifies this bias in the generated videos and provides methods for safe deployment of our models.
+Appendix B Detailed Related Work
 Text-to-Video Generative Modeling.
@@
 This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis.
-7 Conclusion
-We introduced TALC, a simple and effective method for improving the text-to-video (T2V) models for multi-scene generation.
-Specifically, it incorporates the knowledge of the natural alignment between the video segments and the scene-specific descriptions.
-Further, we show that TALC-finetuned T2V model achieve high visual consistency and text adherence while the baselines suffer from one or both of the metrics.
-Given its design, our framework can be easily adapted into any diffusion-based T2V model.
-An important future direction will be to scale the amount of multi-scene video-text data and deploy TALC framework during pretraining of the T2V models.
-8 Acknowledgement
-We would like to thank Ashima Suvarna for providing feedback on the draft. Hritik Bansal is supported in part by AFOSR MURI grant FA9550-22-1-0380.
-Appendix A Prompt for Multi-Scene Caption Generation
+Appendix C Qualitative comparison between baselines and TALC
+We provide an qualitative example to showcase the benefit of TALC over merging captions and merging videos baselines in Figure 6.
+Figure 6: Examples of multi-scene video generation baselines.
+(a) Generating video on the merged descriptions, leads to a poor text-video alignment.
+(b) Generating videos for the individual text descriptions and concatenate them temporally, leads to a lack of visual consistency.
+(c) Our approach (TALC) enhances the scene-level text-video alignment and maintains visual consistency.
+Appendix D Prompt for Multi-Scene Caption Generation
 We present the prompt used to generate multi-scene captions using large multimodal model, Gemini-Pro-Vision, in Figure 7.
@@
 Figure 7: Prompt to generate multi-scene caption using large multimodal models.
-Appendix B Task Prompts
-B.1 Single character in multiple visual contexts
+Appendix E Source datasets for multi-scene video-text data creation
+To construct a multi-scene video-text dataset, we utilize existing dataset that include natural (real) videos and associated high-quality human-written captions that summarize the entire video.
+Specifically, we choose MSR-VTT [53] and VaTeX [54].
+Most of the videos in MSR-VTT are 10-30 seconds long while VaTeX consists 10 seconds long videos.
+In addition, each video in MSR-VTT and VaTex consists 20 captions and 10 captions, respectively, out of which one is selected at random for multi-scene data generation.
+As described above, a single video is cut into multiple video segments using Pyscene library.
+In our experiments, we retain the first four video segments and discard any additional segments if the library generates more than four.
+Since the accuracy of the multi-scene captioning and the computational demands during finetuning are influenced by the number of scenes, we opt to limit the scene count to four for our experiments.
+However, future work could employ similar methodologies to scale the number of scenes, given more computing power and advanced multi-scene captioning models.
+We provide the data statistics for the final multi-scene data in Appendix §K.
+Appendix F Task Prompts
+F.1 Single character in multiple visual contexts
 We present the prompt used to generate multi-scene text descriptions for single character in multiple visual contexts from GPT-4 in Figure 8.
@@
 Figure 8: GPT-4 Prompt to generate multi-scene prompts for single character in multiple visual contexts.
-B.2 Different characters in a specific visual context
+F.2 Different characters in a specific visual context
 We present the prompt used to generate multi-scene text descriptions for different characters in a specific visual context from GPT-4 in Figure 9.
@@
 Figure 9: GPT-4 Prompt to generate multi-scene prompts for different characters in a specific visual context.
-Appendix C Automatic Evaluation Prompt
+Appendix G Automatic Evaluation Prompt
 We present the prompt used to perform automatic evaluation of the multi-scene generated videos using large multimodal model, GPT-4-Vision, in Figure 10.
@@
 Figure 10: Prompt used to perform automatic evaluation of the multi-scene generated videos. We use this prompt when the number of scenes in the task prompt is two.
-Appendix D Human Evaluation Screenshot
+Appendix H Human Evaluation Screenshot
 We present the screenshot for the human evaluation in Figure 11.
@@
 Figure 11: Human Annotation Layout
-Appendix E Fine-grained Analysis
-We present the automatic evaluation results across task prompts (§E.1) and number of scenes (§E.2).
-E.1 Task Prompts
+Appendix I Fine-grained Analysis
+We present the automatic evaluation results across task prompts (§I.1) and number of scenes (§I.2).
+I.1 Task Prompts
 We compare the performance of the baselines and TALC framework across different task prompts in Table 1.
@@
 TALC (F.T.) 93.3 62.6 77.9 98.9 76.3 87.6 79.7 47.9 63.8
-E.2 Number of Scenes
+I.2 Number of Scenes
 We compare the performance of the baselines and TALC framework across different number of generated scenes in Table 2.
@@
 TALC (F.T.) 88.5 66.6 77.5 90.7 64.1 77.4 87.8 56.1 71.9
-Appendix F Inference Details
+Appendix J Inference Details
 We provide the details for sampling multi-scene videos from the ModelScope and Lumiere T2V models in Table 3 and Table 4, respectively.
@@
 Noise scheduler DPMSolverMultiStepScheduler
-Appendix G Multi-Scene Data Statistics
+Appendix K Multi-Scene Data Statistics
 We provide the details for the multi-scene video-text dataset in Table 5.
@@
 Number of video scene-caption instances 20177
-Appendix H Finetuning Details
+Appendix L Finetuning Details
 We provide the details for finetuning ModelScope T2V model with TALC framework in Table 6.
@@
 prediction type epsilon
-Appendix I Single-Scene Video Generation
-Figure 12: Examples of videos generated by Time-Aligned Captions (TALC)
+Appendix M Visual quality of the generated videos.
+Table 7: Human evaluation results on the visual quality of the generated videos from ModelScope.
+We observe that the visual quality of the generated videos are close to each other for the base model.
+However, finetuning the model with merging captions reduces the video quality by a large margin while TALC-finetuned model retains the video quality.
+Method Quality
+Merging Captions (Base) 80.5
+Merging Videos (Base) 86.5
+TALC (Base) 84.5
+Merging Captions (F.T.) 63.4
+TALC (F.T.) 83.3
+We compare the visual quality of the generated videos using human evaluation in Table 7.
+We find that the visual quality of videos generated from the base model ranges from $80.5-86.5$ using the baselines and TALC framework.
+However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points.
+This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video.
+Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining.
+Appendix N Single-Scene Video Generation
 To ascertain that our model’s new multi-scene generation function does not detract from its single-scene generation performance, we conducted a series of evaluations using the VBench framework [56].
@@
 Average 0.63 0.59 0.48
+Figure 12: Examples of videos generated by Time-Aligned Captions (TALC)

### v4

@@
 Abstract
-Recent advances in diffusion-based generative modeling have led to the development of text-to-video (T2V) models that can generate high-quality videos conditioned on a text prompt.
-Most of these T2V models often produce single-scene video clips that depict an entity performing a particular action (e.g., ‘a red panda climbing a tree’).
+Most of these text-to-video (T2V) generative models often produce single-scene video clips that depict an entity performing a particular action (e.g., ‘a red panda climbing a tree’).
 However, it is pertinent to generate multi-scene videos since they are ubiquitous in the real-world (e.g., ‘a red panda climbing a tree’ followed by ‘the red panda sleeps on the top of the tree’).
-To generate multi-scene videos from a pretrained T2V model, we introduce Time-Aligned Captions (TALC) framework.
+To generate multi-scene videos from the pretrained T2V model, we introduce a simple and effective Time-Aligned Captions (TALC) framework.
 Specifically, we enhance the text-conditioning mechanism in the T2V architecture to recognize the temporal alignment between the video scenes and scene descriptions.
-As a result, we show that the pretrained T2V model can generate multi-scene videos that adhere to the multi-scene text descriptions and be visually consistent (e.g., w.r.t entity and background).
-Our TALC-finetuned model outperforms the baseline methods on multi-scene video-text data by 15.5 points on aggregated score, averaging visual consistency and text adherence using human evaluation.
-The project website is https://talc-mst2v.github.io/
-Figure 1: Multi-scene video generation methods.
-(a) Merging scene descriptions; (b) Generating a video per scene independently and concatenating the resulting videos.; (c) TALC: conditioning the generated video on scene-1 description for the first half of the video frames and on scene-2 description for the later video frames.
+For instance, we condition the visual features of the earlier and later scenes of the generated video with the representations of the first scene description (e.g., ‘a red panda climbing a tree’) and second scene description (e.g., ‘the red panda sleeps on the top of the tree’), respectively.
+As a result, we show that the T2V model can generate multi-scene videos that adhere to the multi-scene text descriptions and be visually consistent (e.g., entity and background).
+Further, we finetune the pretrained T2V model with multi-scene video-text data using the TALC framework.
+We show that the TALC-finetuned model outperforms the baseline by achieving a relative gain of $29\%$ in the overall score, which averages visual consistency and text adherence using human evaluation.
+Figure 1: Examples of multi-scene video generation using the TALC framework.
+Our proposed method, TALC, enhances scene-level alignment between text and video, facilitating the generation of videos with seamless transitions between textual descriptions while preserving visual consistency.
+The first two rows are generated using Lumiere [6] and the last two rows use ModelScope [49].
 1 Introduction
 The ability to generate videos that simulate the physical world has been a long-standing goal of artificial intelligence  [1, 47, 10, 36].
-In this regard, text-to-video (T2V) models have seen rapid advancements by pretraining on internet-scale datasets of images, videos, and texts [5, 6, 7, 8, 9].
-Specifically, latest T2V models such as ModelScope [10], Lumiere [11], VideoCrafter [12, 13], and OpenSora [14], train a diffusion model [15] with single-scene video-text datasets [16, 17] that are widely available and easy to acquire.
-As a result, these models can only generate single scene videos that depicts a single action or event [^1] While high-performance T2V models such as Sora [4] might be able to generate multi-scene videos, we point out that they are closed-source models trained with massive compute resources and lack sufficient details on the model design, training protocol, and datasets..
-However, real-world scenarios (e.g., movie and ads creation) often require the generation of multi-scene video for text descriptions (e.g., Scene1: ‘A koala is napping on a tree.’ Scene2: ‘The koala eats leaves on the tree.’).
-In such cases, the generated video should accurately depict the events mentioned in the multi-scene script while maintaining visual consistency (e.g., entity and background consistency).
-Multi-scene T2V generation differs from long video synthesis where the goal is to either interpolate (few frames to many frames) [7] or create continuing patterns of the single event in the generated video [8].
-Prior works [18, 19] use a transformer [20, 21] to generate video frames for a given scene autoregressively.
-However, these models straggle to generate multiple scenes reliably as the context length increases with the history of text descriptions and visual tokens [22] of the previous generated videos, e.g., generating Scene 4 conditioned on Scene1, 2, 3 videos and descriptions.
-Some works [23] utilize a latent diffusion model [24] to generate video frames autoregressively by conditioning on the entire history of generated videos and scene descriptions.
-However, this approach is slow due to repeated sampling, generates only one frame per scene description, and is shown to work with only limited cartoon characters [25, 26] and not with a wide range of visual concepts in the real-world.
-Other works [27, 28] generate multi-scene videos by utilizing knowledge of the entity, background, and their movements from large language models [29].
+In this regard, text-to-video (T2V) models have seen rapid advancements by pretraining on internet-scale datasets of images, videos, and texts [9, 6].
+Previous work [20, 18, 28, 49, 8, 7] primarily focus on training conditional denoising diffusion probabilistic models [22] on paired video-text data [4, 54].
+After training, these models allow for video generation by sampling from the trained diffusion model, conditioned on a text prompt.
+However, most of the open-models such as ModelScope[49] VideoCrafter [13, 14], OpenSora [57] are trained with single-scene video-text dataset [4, 50], which is widely available and easy to acquire.
+However, real-world scenarios often require the generation of multi-scene videos from multi-scene descriptions (e.g., Scene1: ‘A koala is napping on a tree.’ Scene2: ‘The koala eats leaves on the tree.’).
+In such cases, the generated video should accurately depict the events in their temporal order (e.g., Scene2 follows Scene1) while maintaining visual consistency, meaning that backgrounds and entities should remain consistent across scenes.
+While high-performance T2V models such as Sora [36] might be able to generate multi-scene videos, we point out that they are closed-source models trained with massive compute resources and lack sufficient details on the model design, training protocol, and datasets.
+In this work, we present a complementary approach and tackle the challenge of effectively leveraging the capabilities of base T2V models for multi-scene video generation.
+The multi-scene T2V generation differs from long video synthesis where the goal is to either interpolate (few frames to many frames) [18] or create continuing patterns of the single event in the generated video [8].
+Prior work like Phenaki [45, 28] use transformers [44, 3] to generate video frames for a given scene autoregressively.
+However, it is hard for their model to generate multiple scenes reliably as the context length increases with the history of text descriptions and visual tokens [56] of the previous generated videos (e.g., generating Scene 4 conditioned on the Scene1, 2, 3 videos and descriptions).
+Other works [40] utilize a latent diffusion model [41] to generate video frames autoregressively by conditioning on the entire history of generated videos and scene descriptions.
+However, the approach is (a) slow due to repeated sampling, (b) generates only one frame per scene description, and (c) shown to work with only limited cartoon characters [29, 55] instead of wide range of visual concepts in the real-world.
+In this work, our goal is to generate multi-scene videos in the end-to-end manner using a diffusion T2V generative model.
+Prior work like VideoDirectorGPT [30, 31] generates multi-scene videos by utilizing knowledge of the entity, background, and their movements from large language models [2].
 However, these videos are generated independently for each scene before being merged.
-Moreover, these methods do not offer a way to learn from real-world multi-scene video-text data.
-To address these challenges, we propose TALC (Time-ALigned Captions), a simple and effective framework to generate consistent multi-scene videos by effectively leveraging the capabilities of pretrained T2V diffusion models.
-As shown in Figure 1(c), TALC conditions the visual representations of earlier video frames on the embeddings of the earlier scene description, and likewise, it conditions the representations of later video frames on the embeddings of the later scene description in the temporal dimension.
-The temporal modules in the T2V diffusion architecture allows information sharing between video frames (the first half and the second half) to maintain visual consistency.
-Thus, TALC enhances the scene-level text-video alignment while providing all the scene descriptions to the diffusion model at once.
-Further, our TALC framework can enhance the multi-scene text-to-video generation capabilities with real-world multi-scene data (§3.2).
-We provide a qualitative example for comparing different multi-scene video generation methods in Appendix §C.
-In our experiments, we assessed the visual consistency (background and entity consistency) and multi-scene script adherence of the generated videos from Modelscope [10] and Lumiere [11].
-In addition, to our TALC method, we evaluated concatenation approaches, either merging descriptions of different scenes or generating a video per scene independently and concating the output videos (Figure 1(a,b)).
-Through our automatic and human evaluation, we found that merging scene descriptions leads to high visual consistency but poor text adherence.
-On the other hand, merging videos independently achieves high text adherence but visual consistency is compromised.
-Our TALC model strikes an effective balance between visual consistency and text adherence, outperforming the baseline methods by $11.1$ points on the overall score, averaging visual consistency and text adherence scores, as determined by human evaluation.
+To remedy these challenges, we propose TALC (Time-ALigned Captions), a simple and effective framework to generate consistent and faithful multi-scene videos.
+In particular, our approach conditions the T2V generative model with the knowledge of the temporal alignment between the parts of the multi-scene video and multi-scene descriptions (Figure 3).
+Specifically, TALC conditions the visual representations of earlier video frames on the embeddings of the earlier scene description, and likewise, it conditions the representations of later video frames on the embeddings of the later scene description in the temporal dimension.
+Additionally, the temporal modules in the T2V diffusion architecture allows information sharing between video frames (the first half and the second half) to maintain visual consistency.
+Therefore, TALC enhances the scene-level text-video alignment while providing the scene descriptions to the diffusion model all at once (Figure 1).
+Prior methods like FreeNoise [38] propose a motion injection strategy to address these challenges.
+However, this method is quite sophisticated and difficult to control due to diverse hyperparameters, such as time-specific motion injection, prompt interpolation, and injection in specific cross-attention layers.
+In contrast, our approach eliminates these complexities and introduces a straightforward mechanism that significantly boosts performance.
+Unlike previous work, we also demonstrate that finetuning with TALC using real-world multi-scene data can enhance generation capabilities.
+Specifically, we propose a pipeline to curate multi-scene video data and subsequently finetune the T2V model on this multi-scene data using TALC3.2).
+Figure 2: Summary of the results.
+We compare several baselines and our TALC framework with ModelScope [49] for generating multi-scene videos on the unseen prompts.
+Specifically, we study the overall score which averages the visual consistency (object and background) and text adherence (video-text alignment).
+We observe that using TALC with the base model i.e., training-free achieves relative gains of $21\%$ in comparison to FreeNoise [38].
+In addition, we show that multi-scene finetuning with TALC (TALC-FT) allows larger gains, achieving relative gains upto $29\%$.
+In our experiments, we assess the visual consistency (background and entity consistency) and multi-scene script adherence of the generated videos from T2V generative models.
+Through our human evaluation, we find that TALC strikes an effective balance between visual consistency and text adherence, and outperforms the FreeNoise by achieving relative gains of $21\%$ points on the overall score.
+This score represents the average of visual consistency and text adherence scores (Figure 2).
 Furthermore, we construct a multi-scene text-video dataset from real-world videos and fine-tune the T2V generative model using TALC.
-On our human evaluation, the generated videos from the TALC-finetuned model exhibit higher text adherence than the base model in multi-scene scenarios.
-Specifically, it outperforms the baseline methods by $15.5$ points on the overall score.
+We show that finetuning with TALC outperforms FreeNoise by achieving relative gains of $29\%$ on the overall score (Figure 2).
+Further, we perform automatic evaluation for scalable judgements (§5.4) and provide qualitative examples to highlight the benefits of our approach (§5.5).
+We present the related work in Appendix §A.
 2 Preliminaries
-In this work, we focus on generating multi-scene videos from text descriptions using a diffusion-based Text-to-Video (T2V) generative model.
-The initial step is to equip the generative model with the knowledge of a wide range of visual concepts and actions.
-This is achieved during the pretraining stage (§2.1).
-Subsequently, we aim to utilize the pretrained model for multi-scene text-to-video generation task, which we formalize in (§2.3).
 2.1 Diffusion Models for Text-to-Video Generation
 Diffusion models [22, 33] $p_{\theta}(x)$ are a class of generative models that learn data distribution $p_{data}(x)$.
-Their class-conditional versions can be trained to learn conditional distributions $p_{data}(x|y)$ where $y$ is the conditioning variable, which could take various forms such as labels from a dataset or text description accompanying a video [31].
+Due to their flexible design, we can train their class-conditional versions to learn class-conditional data distributions $p_{data}(x|y)$ where $y$ is the conditioning variable.
 We assume a dataset $\mathcal{S}\subset\mathcal{V}\times\mathcal{T}$ consisting of pairs of $(V_{j},T_{j})$ where $V_{j}\in\mathbb{R}^{L\times 3\times H\times W}$ is a raw video consisting of $3$ RGB channels, $L$ frames, $H$ height, $W$ width, and $T_{j}$ is a text caption.
-The T2V model is trained to learn the conditional distribution of the videos conditioned on the text $p_{\mathcal{S}}(V_{j}|T_{j})$.
-In this work, we consider diffusion-based generative models that learn this distribution via iterative denoising of the video $z_{j}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$.
-Here, $z_{j}$ can either represent the video in the raw pixel space $V_{j}$ [11] or it can represent the latent representation of the video $z_{j}=\mathcal{E}(V_{j})$ for the latent diffusion models [24] where $\mathcal{E}$ is an encoder network such as VAE [32].
-Given $z_{j}$, diffused variable $z_{\tau,j}=\alpha_{\tau}z_{j}+\beta_{\tau}\epsilon$ are constructed where $\epsilon\sim\mathcal{N}(0,I)$ and $\alpha_{\tau}$ and $\beta_{\tau}$ are sampled from the noise scheduler $p_{\tau}$ [33], which defines the noise levels the model is trained on.
-Finally, a denoiser network $f_{\theta}$ [34, 35], which takes the diffused variable $z_{\tau}$ and embeddings of the text caption as input, is trained to predict the target vector $y$ where $y$ can be the original noise $\epsilon$, which minimizes the denoising score matching objective [15]:
+We use $\mathcal{V}$ and $\mathcal{T}$ to denote the domain of videos and text, respectively.
+The aim of T2V generative modeling is to learn the conditional distribution of the videos conditioned on the text $p_{\mathcal{S}}(V_{j}|T_{j})$.
+In this work, we consider diffusion-based generative models that learn the data distribution via iterative denoising of the input video $z_{j}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$.
+Here, $z_{j}$ can either represent the input video in the raw pixel space $V_{j}$ [6] or it can represent the latent representation of the video $z_{j}=\mathcal{E}(V_{j})$ for the latent diffusion models [41] where $\mathcal{E}$ is an encoder network [27].
+Given $z_{j}$, diffused variable $z_{\tau,j}=\alpha_{\tau}z_{j}+\beta_{\tau}\epsilon$ are constructed where $\epsilon\sim\mathcal{N}(0,I)$ where $\alpha_{\tau}$ and $\beta_{\tau}$ are sampled from the noise scheduler $p_{\tau}$ [16].
+Finally, we train a denoiser network $f_{\theta}$ [42, 37] that inputs the diffused variable $z_{\tau}$ and embeddings of the text caption to predict the target vector $y$ where $y$ can be the original noise $\epsilon$, which minimizes the denoising objective [22]:
 $\mathbb{E}_{(V_{j},T_{j})\in S,\tau\sim p_{\tau},\epsilon\sim N(0,I)}\left[||\epsilon-f_{\theta}(\tau,z_{\tau,j},h_{j})||_{2}^{2}\right]$ (1)
-where $h_{j}=\mathcal{H}(T_{j})\in\mathbb{R}^{d}$ is the embedding of the text caption $T_{j}$, $\mathcal{H}$ is the text embedding model [36] and $d$ is the dimension size.
+where $h_{j}=\mathcal{H}(T_{j})\in\mathbb{R}^{d}$ is the embedding of the text caption $T_{j}$ where $\mathcal{H}$ is the text embedding model [39] and $d$ is the dimension size.
 2.2 Text Conditioning Mechanism
-For effective textual controllability of video generation, the denoiser networks is equipped with a cross-attention mechanism [10, 7].
+To ensure the effective textual controllability of video generation, the structure of the denoiser networks is equipped with a cross-attention mechanism [49, 18].
 Specifically, it conditions the visual content $z_{\tau}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ on the text.
@@
 Intuitively, the repeat operation represents that the $L$ frames of the video $z_{j}$ are semantically aligned with the textual description $T_{j}$ or its text embedding $r_{j}$.
-In §3, we will manipulate this operation to make the model architecture aware of the video-text alignment in the multi-scene scenario.
 These repeated text embeddings $r_{j}$ are inputs to the spatial attention block as the key and value in the multi-head attention block.
-The cross-attention enables the intermediate visual features to capture the semantic information that facilitates an alignment between the text and vision embeddings.
-Formally,
+The cross-attention enables the intermediate visual features to capture the semantic information that facilitates an alignment between the language and vision embeddings.
 $z^{\prime}_{\tau,j}=CA_{f_{\theta}}(Q=z_{\tau,j};K=r_{j};V=r_{j})$ (2)
-where $CA_{f_{\theta}}$ is the cross attention mechanism with $Q,K,V$ as the query, key, and value, respectively, in the spatial blocks of the denoiser network, $z^{\prime}_{\tau,j}$ is the intermediate representation that is informed with the visual and textual content of the data.
-In addition to the spatial blocks, the denoiser network also consists of temporal blocks that aggregate features across video frames, which are useful for maintaining visual consistency in the generated video.
+where $CA_{f_{\theta}}$ is the cross attention mechanism with $Q,K,V$ as the query, key, and value, respectively, in the spatial blocks of the denoiser network.
+Additionally, $z^{\prime}_{\tau,j}$ is the intermediate representation that is informed with the visual and textual content of the data.
+In addition to the spatial blocks, the denoiser network also consists temporal blocks that aggregate features across video frames which are useful for maintaining visual consistency in the video.
 2.3 Multi-Scene Text-to-Video Generation
-We aim to generate multi-scene video $X=\{x_{1},x_{2},\ldots,x_{n}\}$ from multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ where $n$ are the number of scenes and each sentence $y_{j}$ is a description for scene $j$.
-Index $j$ also defines the temporal order of events in the multi-scene script, i.e., if $k>j$, we want the events described in scene $j$ to be depicted earlier than the events described in scene $k$.
-Further, we want that each video part $x_{j}$, within the entire generated video $X$, to have high video-text semantic alignment with its corresponding scene description $y_{j}$.
-Figure 2: Time-Aligned Captions (TALC) Framework.
-During the generation process of the video, the initial half of the video frames are conditioned on the embeddings of the description of scene 1, while the subsequent video frames are conditioned on the embeddings of the description of scene 2.
+In this work, we aim to generate multi-scene video $X=\{x_{1},x_{2},\ldots,x_{n}\}$ from multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ where $n$ are the number of sentences and each sentence $y_{j}$ is a scene description for scene $j$.
+Additionally, the index $j$ also defines the temporal order of events in the multi-scene script i.e., we want the events described in the scene $j$ to be depicted earlier than the events described in the scene $k$ where $k>j$.
+Further, we want the parts of the entire generated video $X$, given by $x_{j}$, to have high video-text semantic alignment with the corresponding scene description $y_{j}$.
+In addition, we expect the appearance of the objects to remain consistent throughout the video unless a change is specified in the description.
 3 TALC: Time-Aligned Captions for Multi-Scene T2V Generation
-In this section, we present our novel method TALC, describing its architecture and a novel approach for constructing synthetic multi-scene video-text data to train TALC.
-3.1 TALC Architecture
-Our starting point is that each specific scene $x_{j}$ in a multi-scene video $X\in\mathbb{R}^{L\times 3\times H\times W}$ should reflect the entities and events described in the conditioning scene description $y_{j}$.
-To achieve this, TALC adapts T2V diffusion architecture $f_{\theta}$ to multi-scene generation by ensuring that the representation for video part $x_{j}$ attends and aggregates language features specifically from description $y_{j}$.
-Concretely, TALC allocates $L/n$ frames within the entire video to each $x_{j}\in X$.
-Let $z_{X}=[z_{x_{1}};z_{x_{2}};\ldots;z_{x_{n}}]\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ be the representation for the entire video $X$, and $z_{x_{j}}\in\mathbb{R}^{(L/n)\times C\times H^{\prime}\times W^{\prime}}$ for the different scenes, which are concatenated in the temporal dimension.
-In addition, $r_{Y}=\{r_{y_{1}},\ldots,r_{y_{n}}\}$ be the sequence of text embeddings for the multi-scene description $Y$ and $y_{j}$ be an individual scene description.
-Under TALC, Eq.
+3.1 Approach
+Most of the existing T2V generative models [49, 13, 6] are trained with large-scale short video-text datasets (10 seconds - 30 seconds) such as WebVid-10M [4].
+Here, each instance of the dataset consists of a video and a human-written video description.
+These videos either lack the depiction of multiple events, or the video descriptions do not cover the broad set of events in the video, instead focusing on the major event shown.
+As a result, the pretrained T2V generative models only synthesize single video scenes depicting individual events.
+We introduce TALC, a novel and effective framework to generate multi-scene videos from diffusion T2V generative models based on the scene descriptions.
+Our approach focuses on the role of text conditioning mechanism that is widely used in the modern T2V generative models (§2.2).
+Specifically, we take inspiration from the fact that the parts of the generated video $x_{j}$ should depict the events described in the scene description $y_{j}$.
+To achieve this, we ensure that the representations for the part of the generated video aggregates language features from the scene description $y_{j}$.
+Consider that we want to generate a multi-scene video $X\in\mathbb{R}^{L\times 3\times H\times W}$ from the scene descriptions $y_{j}\in Y$, using a T2V generative model $f_{\theta}$.
+Furthermore, we assume that individual video segments $x_{j}$ are allocated $L/n$ frames within the entire video $X$.
+Let $z_{X}=[z_{x_{1}};z_{x_{2}};\ldots;z_{x_{n}}]\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ represent the representation for the entire video $X$, and $z_{x_{j}}\in\mathbb{R}^{(L/n)\times C\times H^{\prime}\times W^{\prime}}$ for the $j^{th}$ part of the video that are concatenated in the temporal dimension.
+In addition, consider $r_{Y}=\{r_{y_{1}},\ldots,r_{y_{n}}\}$ be the set of text embeddings for the multi-scene description $Y$ and $y_{j}$ be an individual scene description.
+In the TALC framework, the Eq.
 2 is changed to:
@@
 $z^{\prime}_{\tau,X}$ $=[z^{\prime}_{x_{1}};z^{\prime}_{x_{2}};\ldots;z^{\prime}_{x_{n}}]$ (4)
-with, $\tau$ the timestamp in the diffusion modeling setup, which is applied during training as well as inference. We illustrate the framework in Figure 2.
-While TALC aims to equip the generative model with the ability to follow all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the existing denoiser network.
-By design, our approach can be applied to the pretrained T2V model during inference.
-3.2 TALC Training
-Figure 3: Our approach for generating time-aligned video captions.
-The process begins with PyScene cuts identifying the boundaries of distinct scenes within a video.
-Keyframes are then selected from the median of each scene.
-These frames are processed collectively through the Gemini model to produce multi-image captions that maintain narrative continuity by contextualizing each scene within the video’s overall sequence.
-We note that the text adherence capabilities of the pretrained T2V generative model are limited.
-This is due to the lack of multi-scene video-text data during pretraining, since these datasets are not widely available and are hard to curate for model training.
-Prior work, such as ActivityNet [37], curated human captions for specific video scenes depicting useful actions in long videos.
-However, these scenes are either overlapping or have a large temporal gap between them, which are less suitable for natural and smooth transitions required in multi-scene videos.
-To this end, we propose a method for creating synthetic real-world multi-scene video-text examples, with which we further train our T2V model.
-Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description.
-Then, similar to [9], we apply PySceneDetect [^2] https://github.com/Breakthrough/PySceneDetect to segment $A_{i}$ into continuous video scenes $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of detected scenes.
-Next, we sample the middle video frame $F_{i,j}$ in each scene $A_{i,j}$ as its semantic representative.
-Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for $A_{i}$ and the entire video caption $B_{i}$ to Gemini-Pro-Vision-1.0 [38].
-Gemini is prompted to generate captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$.
-In Figure 3, we provide an instance for the multi-scene video-text data generation.
-We provide the prompt provided to Gemini and the source datasets in Appendix §D and §E, respectively.
-4 Experimental Setting
-In this section, we describe the setup of our experiments for generated multi-scene videos, including eval-set creation, metrics, automatic and human evaluation, and the tested models.
-4.1 Eval-Set Task Prompts
-To create an eval-set for testing models that generate multi-scene videos, we curate a set of task prompts for diverse scenarios (The full task prompts are in Appendix §F.).
-We generate $100$ prompts for each scenario.
-Single character in multiple visual contexts (S1): We instruct GPT-4 [39] to create a coherent script consisting of four scenes.
+Here, $\tau$ represents the timestamp in the diffusion modeling setup, which is applied during training as well as inference.
+We illustrate the framework in Figure 3.
+While TALC aims to equip the generative model with the ability to depict all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the denoiser network.
+By design, our approach can be applied to the base T2V model without any further training.
+Figure 3: The architecture of Time-Aligned Captions (TALC).
+During the generation process of the video, the initial half of the video frames are conditioned on the embeddings of the description of scene 1 ($r_{y_{1}}$), while the subsequent video frames are conditioned on the embeddings of the description of scene 2 ($r_{y_{2}}$).
+3.2 Multi-Scene Video-Text Data Generation
+While our approach generates better multi-scene videos, the text adherence capabilities of the pretrained T2V generative model are limited.
+This is due to the lack of multi-scene video-text data during its pretraining.
+To this end, we create a real-world multi-scene video-text dataset to allow further training of the pretrained T2V models.
+Specifically, we leverage the capability of the multimodal foundation model, Gemini-Pro-Vision [43], to generate high-quality synthetic data for enhanced video-text training [5].
+Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description from the dataset.
+Subsequently, we utilize PySceneDetect library  [^1] https://github.com/Breakthrough/PySceneDetect to generate continuous video scenes from $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of scene cuts in the video.
+Further, we sample the middle video frame $F_{i,j}$ as a representative of the semantic content in the video scene $A_{i,j}$.
+Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for a single video $A_{i}$ and the entire video caption $B_{i}$ to Gemini-Pro-Vision.
+Specifically, the model is prompted to generate high-quality captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$.
+We provide the prompt provided to the multimodal model in Appendix §E.
+In Figure 4 we provide an instance for the multi-scene video-text data generation.
+We highlight that higher-quality multi-scene datasets would enhance the performance of the models with TALC framework.
+We provide the more details in Appendix C.
+[^2] We note that the parts of this pipeline are also utilized in a concurrent work, ShareGPTVideo [15].
+However, their approach is focused on long-form video generation pretraining, while we are focus on multi-scene video generation for finetuning.
+4 Evaluation
+4.1 Metrics
+The ability to assess the quality of the generated multi-scene videos is a challenging task itself.
+As humans, we can judge the multi-scene videos across diverse perceptual dimensions [25] that the existing automatic methods often fails to capture [11].
+Following [30], we focus on the visual consistency of the generated video, and text adherence capabilities of the T2V models.
+Visual Consistency.
+This metric aims to assess the (entity or background) consistency between the frames of the multi-scene videos.
+Here, the entity consistency aims to test whether the entities in the multi-scene video are consistent across the video frames.
+For instance, the appearance of an animal should not change without a change described in the text description.
+In addition, the background consistency aims to test whether the background of the multi-scene video remains consistent across the video frames.
+For instance, the room should not change without a change in text description.
+Text Adherence.
+This metric aims to test whether the generated video adheres to the multi-scene text description.
+For instance, the events and actions described in the text script should be presented in the video accurately, and in the correct temporal order.
+In our experiments, we compute the visual consistency and text adherence with the human and automatic evaluators.
+Further, we compute the overall score, which is the average of the visual consistency and text adherence scores.
+In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Appendix D).
+4.2 Task Prompts
+Here, we curate task prompts for diverse scenarios to holistically assess the quality of the videos.
+Single character in multiple visual contexts (S1).
+In this scenario, we instruct an LLM, GPT-4 [2], to create a coherent script consisting of four scenes.
 Each scene features a specific animal character performing diverse activities in every scene.
-This task assesses the consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script.
-An example script is “Scene 1: A red panda is climbing a tree.
-Scene 2: The red panda eats the leaves on the tree.
-Scene 3: … Scene 4: …”.
-Different characters in a specific visual context (S2): We instruct GPT-4 to create a coherent script consisting of four scenes.
-Each scene features the same activity but with different animal characters [18].
-This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the animal character changes.
-An example script is “Scene 1: A cat leaps onto a countertop.
-Scene 2: A dog leaps onto the same countertop.
-Scene 3: A rabbit … Scene 4: A raccoon …”.
-Multi-scene captions from real videos (S3): We use our multi-scene video/text data generation pipeline (§3.2) to create similar prompts for real videos for the test splits of the video-text datasets.
-This task assesses the ability of a model to generate multi-scene videos with consistent appearances of entities and background for open-ended prompts, which are derived from real videos.
-An example script is “Scene 1: A vlogger introduces her skincare routine.
-Scene 2: She applies a serum to her face.
-Scene 3: …”.
-4.2 Evaluation Procedure
-Following [27], we assess the capability of the tested T2V models to: (a) follow the text instructions in the script (Text Adherence); (b) keep consistent appearance of entities and background across scenes (Visual Consistency), unless the script specifically requests a change in any of them.
-We compute the visual consistency and text adherence with automatic and human evaluators.
-We also report an overall score, which is the average of the visual consistency and text adherence scores.
-In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Appendix M):
-Automatic Evaluation: We utilize GPT-4-Vision (GPT4V) [40] to reason over multiple images.
-First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for a two-scene video).
-Then, we prompt GPT4V with the temporal sequence of the sampled frames and the multi-scene text description and instruct GPT4V to decide on the quality of the generated video across several attributes, including entity consistency, background consistency, and text adherence.
-For each attribute, GPT4V assigns one of three possible response $\{\textit{yes}=1,\textit{partial}=0.5,\textit{no}=0\}$.
-The automatic evaluation prompt is given in Appendix §G.
-[…]
-Since merging captions (§4.3.1) and TALC3.1) methods input the entire multi-scene text description at once, the quality of the video generated by these methods is influenced by the number of scenes described in the text script.
-Hence, we calculate the performance of the tested models by averaging the scores assigned to videos generated for two, three, and four scenes, over 900900900900333100 (examples per scenario) * 3 (scenarios) * 4 (2-scene/3-scene/4-scene) = 900.
-multi-scene prompts.
-We report visual consistency by averaging the performance across the entity and background consistency metrics.
-Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), while the background consistency and text adherence are computed for all the task prompts.
-4.3 Tested Models
-We experiment with ModelScope [10] and Lumiere [11] T2V models for multi-scene video generation with TALC.
-We mainly experiment with ModelScope due to its ease of reproduction and adoption in prior work [27].
-We also experiment we Lumiere to showcase TALC’s generalization.
-By design, TALC can be applied to a base T2V model during inference, without any multi-scene finetuning.
-We compare the multi-scene performance of the two base models without finetuning under three settings: TALC, merging captions, and merging videos (see baseline details in Section 4.3.1).
-In addition, we finetune ModelScope with the TALC framework using the multi-scene video-text data (§3.2).
-As a baseline, we also finetune ModelScope without the TALC framework by naively merging the scene-specific captions in the raw text space.
-In both cases, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$.
-We provide more finetuning details in Appendix §L.
-During inference, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere. More inference details are found in Appendix §J.
-4.3.1 Baselines
-For Merging Captions, we create a single caption by merging all the multi-scene descriptions.
-Specifically, the multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ can be written as a single prompt ‘$\mathcal{P}=y_{1}.\text{Then, }y_{2}.\ldots\text{Then, }y_{n}.$’.
-For instance, the two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’} would change to $\mathcal{P}=$A red panda climbs on a bamboo forest.
-Then, the red panda sleeps peacefully in the treetop.” Subsequently, we generate a video from the T2V model $f_{\theta}$ by conditioning it on $\mathcal{P}$.
-For Merging Videos, we generate a video $x_{i}$ for each scene description $y_{i}$ individually and stitch the output videos together afterwards to synthesize the entire video $X={x_{1},x_{2},\ldots,x_{n}}$.
+This task assesses the capability of the T2V model to generate consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script.
+Different characters in a specific visual context (S2).
+In this scenario, we instruct a language model, GPT-4, to create a coherent script consisting of four scenes.
+Each scene features different animal characters engaging in the same activity in every scene [45].
+This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the appearance of the different characters in the multi-scene text script.
+Multi-scene captions from real videos (S3).
+Here, we aim to assess the ability of the model to generate multi-scene videos for open-ended prompts that are derived from real-world videos.
+This task also assesses the ability of the T2V model to generate consistent appearances of the entity and its background while adhering to multi-scene descriptions.
+Specifically, we use our multi-scene video-text data generation pipeline (§3.2) to create such prompts for the real videos from the test splits of the video-text datasets.
+In total, we generate $100$ prompts in this scenario.
+We present a sample of the various task prompts in the Appendix §F.
+We present the details about the human and automatic evaluators in §B.
+4.3 Evaluation Setup
+We compute the performance of the baselines and TALC by averaging the scores assigned to videos generated for two, three, and four scenes.
+Additionally, we report on visual consistency by averaging the performance across the entity and background consistency metrics.
+Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), and the background consistency and text adherence scores are computed for all the task prompts.
+We also evaluate the impact of TALC-based finetuning for single-scenes in Appendix §M.
+Table 1: Human evaluation results.
+We present the human evaluation results for the overall score for several baselines (e.g., FreeNoise, merge videos, and merge captions) and TALC framework for the ModelScope generative model.
+Specifically, we find that the TALC with the base model outperforms the FreeNoise by achieving the relative gain of $21\%$ on the overall score.
+In addition, finetuning ModelScope with TALC framework enables better multi-scene video generation by achieving the highest overall score.
+Overall, TALC strikes a good balance between visual consistency and text adherence.
+Overall score Visual consistency Text adherence
+FreeNoise [38] 59.7 77.0 42.5
+Merge videos 61.3 ($+2.6\%$) 55.0 67.5
+Merge captions 64.8 ($+8.5\%$) 96.5 33.0
+Finetuning w/ Merge captions 61.1 ($+2.3\%$) 80.0 42.3
+TALC 72.4 ($+21\%$) 92.3 52.5
+Finetuning w/ TALC 76.8 ($+29\%$) 86.4 67.2
 5 Experiments
-(a) Performance on ModelScope T2V model.
-(b) Performance on Lumiere T2V model.
-Figure 4: Automatic evaluation results for (a) ModelScope and (b) Lumiere.
-In (a), we observe that TALC-finetuned ModelScope model achieves the highest overall score, that is the average of the visual consistency and text adherence scores.
-In (b), we find that TALC framework with the Lumiere base model outperforms merging captions and merging videos on the overall scores.
-We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
-5.1 Automatic Evaluation Results
-We compare the performance of the baselines (e.g., merging captions and merging videos) with the TALC framework for ModelScope and Lumiere using the automatic evaluation in Figure 4.
+5.1 Text-to-Video Generative Models
+We perform most of our experiments on ModelScope [49] due to its easy-of-access and adoption in prior works [30].
+In addition, we also include Lumiere-T2V, a model that leverages space-time U-Net denoising networks to generate high-quality videos.
+In this work, we include early experiments with Lumiere to showcase the flexibility of the TALC approach on diverse models.
+We perform human evaluation for ModelScope, and automatic evaluation for both ModelScope and Lumiere.
+Base model with TALC.
+As described in §3.1, our approach modifies the traditional text-conditioning mechanism to be aware of the alignment between text descriptions and individual video scenes.
+By design, the TALC framework can be applied to the base T2V model during inference, without any multi-scene finetuning.
+Here, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere.
+We provide more details in Appendix §J.
+Finetuning with TALC.
+Since the base model is pretrained with single-scene data, we aim to show the usefulness of TALC framework when we have access to the multi-scene video-text data.
+To this end, we finetune ModelScope on the multi-scene video-text data (§3.2) with TALC framework.
+As a pertinent baseline, we also finetune the ModelScope without TALC framework by naively merging the scene-specific captions in the raw text space.
+In this setting, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$.
+We provide further details on the finetuning setup in Appendix §L.
+The inference settings are identical to the prior method of generating videos from the base model without finetuning.
+In this section, we present the results for the baselines and TALC framework averaged over a diverse task prompts and multiple scenes using automatic evaluation (§5.4) and human evaluation (§5.3).
+Finally, we provide qualitative examples for the multi-scene generated videos to showcase the usefulness of our approach (§5.5).
+5.2 Baselines
+Merge Captions
+In this setup, we create a single caption by merging all the multi-scene descriptions.
+While this approach mentions the temporal sequence of the events in a single prompt, the T2V model does not understand the temporal boundaries between the two events.
+Specifically, the visual features for all the frames will aggregate information from the entire multi-scene description, at once, without any knowledge about the alignment between the scene description and its expected appearance.
+Merge Videos
+In this setup, we generate videos for each scene description individually and merge them in the raw input space.
+In this process, the parts of the multi-scene video closely adhere to the scene descriptions, leading to high text fidelity.
+However, since the generated videos do not have access to all the multi-scene descriptions (e.g., the video for Scene 2 is not informed about Scene 1), the visual consistency across the entire video is quite poor.
+FreeNoise
+We also consider a more sophisticated approach, FreeNoise [38].
+Specifically, it reschedules a sequence of noises for long-range correlation and perform temporal attention over them by window-based fusion.
+In addition, it includes motion injection method to support the generation of videos conditioned on multiple text prompts.
+Due to its complexity, this method involves setting various hyperparameters, including the denoising step at which motion injection should be activated, the cross-attention layers where prompts are injected, and the frames between which the prompt representations should be interpolated.
+In this work, we evaluate the publicly available implementation with VideoCrafter [13] on our unseen prompts.
+[^3] https://github.com/AILab-CVC/FreeNoise While Gen-L-Video [48] has been shown to perform worse than FreeNoise, we conduct a qualitative analysis to emphasize the robustness of our method against it.
+(§5.5).
+[^4] Other approaches such as VideoDirectorGPT [30] and Phenaki [45] are not publicly available.
+5.3 Human Evaluation
+TALC achieves the best performance in human evaluation.
+We compare the performance of the baselines and TALC framework using human evaluation in Table 1.
+We find that TALC applied to ModelScope (MS) outperforms FreeNoise by achieving a relative gain of $21\%$ on the overall score.
+In addition, finetuning MS with TALC framework increases the video generation capability, leading to relative gain of $27\%$ over FreeNoise in the overall score.
+The poor performance of FreeNoise can be attributed to the inability to introduce new content and limited motions as the multiple scenes ($>2$) are requested.
+We provide some qualitative examples for FreeNoise to highlight its limitations in Appendix N.
+In addition, we find that using TALC framework in the base model outperforms the merging captions and merging video methods with the base model by $7.6$ points and $11.1$ points, respectively, on the overall score.
+Further, we note that TALC-finetuned model outperforms the merging captions and merging video methods with the base model by $12$ points and $15.5$ points, respectively, on the overall score.
+Additionally, we observe that the merging captions with the base model achieves the highest visual consistency score of $96.5$ points while it is the lowest for merging videos generated from the base model.
+Our results indicate that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames.
+Further, we note that the text adherence of the TALC-finetuned and TALC-base model is better than merging captions-finetuned and merging captions-base model, respectively.
+The high text adherence for merging videos can be attributed to its design where individual video scenes adhere to the scene-specific descriptions well.
+Overall, our empirical findings highlight that our simple framework can enable robust multi-scene video generation with finetuning.
+We present the results for the visual quality metric in Appendix D.
+Table 2: Automatic evaluation results for ModelScope.
+We compare the performance of the baselines and TALC framework using the automatic evaluation for ModelScope generative model.
+Similar to the human evaluation, we observe that our simple approach achieves the best overall score with the base model (TALC) and finetuned model (Finetuning w/ TALC).
+We abbreviate visual consistency as VC, and text adherence as TA.
+Overall score VC TA
+Merge captions 61.7 91.0 32.4
+Merge videos 67.5 ($+9.4\%$) 65.0 70.0
+Finetuning w/ Merge captions 57.3 ($-7.1\%$) 77.0 37.5
+TALC 68.6 (+11.2%) 89.9 47.2
+Finetuning w/ TALC 75.6 (+22.5%) 89.0 62.3
+Table 3: Automatic evaluation results for Lumiere.
+We present the results for the comparison between the baselines and TALC for the Lumiere video generative model.
+Specifically, our results indicate that TALC achieves the highest overall score, by achieving $7.1\%$ relative gains over merge captions on the overall score.
+This indicates that TALC is a flexible strategy that can be applied to diverse video generative models.
+We abbreviate visual consistency as VC, and text adherence as TA.
+Overall score VC TA
+Merge captions 64.4 94.7 34
+Merge videos 66.5 (+3.2%) 68.0 65.0
+TALC 69.0 (+7.1%) 97.8 40.0
+5.4 Automatic Evaluation
+We compare the performance of the baselines with the TALC framework for ModelScope and Lumiere using the automatic evaluation in Table 2 and 3, respectively.
 TALC outperforms the baselines without any finetuning.
-In Figure 4(a), we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($69$ points), outperforms the overall score achieved by the videos generated using merging captions ($62$ points) and merging videos ($68$ points) with the base ModelScope.
-Specifically, we observe that the visual consistency of the generated video is high for merging captions ($91$ points) and TALC ($89.9$ points) while it is low for merging videos ($65$ points).
-This indicates that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames.
-Since merging captions mentions multiple events in a single prompt, the T2V model struggles to understand the temporal boundaries between these events.
-As indicated by Eq.
-2, the visual features for all video frames aggregate information from the entire multi-scene description simultaneously, without aligning individual scene descriptions with their corresponding appearances in the generated video.
-In Figure 4(a), this results in poor text adherence with merging captions ($32$ points).
+In Table 2, we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($68.6$ points), outperforms the overall score achieved by the videos generated using merging captions ($61.7$ points) and merging videos ($67.5$ points) with the base ModelScope.
 In addition, we observe that the text adherence using TALC outperforms merging captions by $14.8$ points, while the text adherence is the highest with a score of $70$ points using merging videos.
-This can be attributed to the design of the merging videos baseline where individual video scenes adhere to the scene-specific descriptions well.
-Hence, merging videos independently approach can be viewed as an upper bound on the text adherence metric.
-In Figure 4(b), we observe similar trends for the Lumiere T2V generative model.
+In Table 3, we observe similar trends for the Lumiere T2V generative model.
 Specifically, we find that the overall score for TALC outperforms merging captions and merging videos by $4$ points and $2$ points, respectively.
@@
 Further, we find that TALC outperforms merging captions by $5$ points on text adherence, while merging videos achieves the highest text adherence $65$ points.
-This highlights that the TALC framework enhances the adherence to the scene descriptions.
+This highlights that the model more easily generates multi-scene videos that adhere to individual text scripts, whereas adherence to the text diminishes when the model is given descriptions of multiple scenes all at once.
 Finetuning with TALC achieves the best performance.
-Earlier, we evaluated the usefulness of the TALC framework with the base model.
-However, the base models are trained with the single-scene video-text data that might limit their capability for multi-scene video generation.
-To alleviate this issue, we finetune ModelScope T2V model on the multi-scene video-text data (§3.2).
-Specifically, we finetune the model using the merging captions method and TALC framework, independently.
-In Figure 4(a), we find that finetuning with TALC achieves the highest overall score of $75.6$ points in comparison to all the baselines.
+In Table 2, we find that finetuning with TALC achieves the highest overall score of $75.6$ points in comparison to all the baselines.
 Specifically, we observe that the visual consistency does not change much with finetuning using the TALC method ($89.9$ points vs $89$ points).
@@
 Additionally, we find that the text adherence of the TALC-finetuned model is $15.1$ points more than the text adherence of the TALC-base model.
-Similarly, we find that the text adherence of the merging captions-finetuned model is $5.1$ points more than the text adherence of the merging captions-base model.
-This highlights that finetuning a T2V model with multi-scene video-text data helps the most with enhancing its text adherence capability.
+This highlights that finetuning a T2V model with multi-scene data helps the most with its text adherence capability.
 Fine-grained Results.
@@
 Overall, we show that the TALC strikes a good balance between visual consistency and text adherence to generate high-quality multi-scene videos.
-5.2 Human Evaluation Results
-Figure 5: Human evaluation results for ModelScope model.
-We observe that the base model using the TALC framework outperforms the merging captions and merging videos baselines on the overall score.
-In addition, TALC-finetuned model enhances the text adherence and achieves the highest overall score.
-We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
-TALC achieves the best performance in human evaluation.
-We compare the performance of the baselines and TALC framework for ModelScope using human evaluation in Figure 5.
-We find that TALC-finetuned model outperforms the merging captions and merging video methods with the base model by $12$ points and $15.5$ points, respectively, on the overall score.
-In addition, we find that using TALC framework in the base model outperforms the merging captions and merging video methods with the base model by $7.6$ points and $11.1$ points, respectively, on the overall score.
-Further, we observe that the merging captions with the base model achieves the highest visual consistency score of $96.5$ points while it is the lowest for merging videos generated from the base model.
-In addition, we find that the text adherence of the TALC-finetuned and TALC-base model is better than merging captions-finetuned and merging captions-base model, respectively.
-Our results highlight at the benefit of including the inductive bias of temporal alignment between the video scenes and their scene descriptions for multi-scene video generation.
-We perform human evaluation on the visual quality of the generated multi-scene videos in §M.
-While our work is centered around multi-scene evaluation, we also perform single-scene evaluation in Appendix §N.
 5.5 Qualitative Analysis
-We provide qualitative examples of generating multi-scene videos using ModelScope (fine-tuned with TALC) and Lumiere (base model with TALC) for diverse scenarios in Figure 12.
-Our analysis reveals that both ModelScope and Lumiere are capable of producing multi-scene videos that exhibit high text adherence and visual consistency.
-Considering the case of the same animal engaging in multiple actions (referred to as ‘one character multiple contexts’).
-The videos generated by ModelScope successfully maintained the same animal while varying the background and action between the scenes.
-Conversely, the videos generated by Lumiere displayed the same animal performing different actions with minimal background alterations.
-We believe that this distinction is attributed to ModelScope’s fine-tuning with TALC.
-Considering different animals within a particular visual setting (referred to as ‘multiple-characters same context’), both ModelScope and Lumiere demonstrated impressive abilities in preserving the consistency of the background across the videos and adhering closely to the provided text.
-During our analysis, we noticed that the multi-scene captions derived from real videos (referred to as ‘open-ended captions’) exhibited a substantial number of changes between the various scenes.
-In this scenario, Lumiere, when employed without fine-tuning, displayed challenges in adhering to the text, while ModelScope achieved a higher degree of text adherence but was also prone to visual artifacts.
-6 Related Work
-Text-to-Video Generative Modeling.
-The field of text-to-video (T2V) synthesis has significantly evolved from its inception with models like VGAN [2] and MoCoGAN [41], leveraging the foundational technologies of GANs [42] and VAEs [43] to produce concise, single-scene videos.
-The advent of diffusion models, exemplified by Imagen Video [44], marked a notable advancement in T2V synthesis.
-Despite these strides, the challenge of creating multi-scene videos that reflect the complexity of the physical world [1, 2, 3] remains.
-Our work, TALC, extends the capabilities of T2V models to multi-scene storytelling, filling a crucial gap in the synthesis landscape.
-Multi-Scene Video Generation.
-The pursuit of multi-scene T2V synthesis has been furthered by recent innovations like Phenaki [18] and Stable Video Diffusion [9], which have explored new frontiers in video generation from textual prompts and the scaling of latent diffusion models, respectively.
-Despite these advancements, the task of generating videos that convey coherent narratives across multiple scenes remains formidable, highlighted by recent works such as VideoPoet [19].
-TALC tackles this task and offers a framework produces videos spanning multiple scenes.
-We also introduce nuanced evaluation approach.
-This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis.
-We provide a detailed related work in §B.
+We provide qualitative examples of generating multi-scene videos using TALC, FreeNoise and Gen-L-Video in Appendix Figure 10 and Figure 11.
+Our analysis demonstrates that all methods are capable of generating multi-scene videos that exhibit a high degree of text adherence.
+However, the primary distinction lies in the quality of the videos.
+FreeNoise is capable of producing videos characterized by superior quality and visual coherence.
+Nonetheless, the motion within these videos is notably limited, resulting in a relatively static scene.
+Gen-L-Video can generate videos that incorporate motion.
+However, visual consistency is not maintained throughout the video.
+For instance, in the provided example depicting a man engaged in both surfing and skiing, the man’s appearance undergoes noticeable changes across the video.
+Overall, we observe that TALC is capable of generating realistic multi-scene videos that not only exhibit high textual adherence but also maintain visual consistency.
+Furthermore, the transitions between scenes are smooth and natural.
 6 Conclusion
@@
 We would like to thank Ashima Suvarna for providing feedback on the draft. Hritik Bansal is supported in part by AFOSR MURI grant FA9550-22-1-0380.
-Appendix A Limitations
-While there are numerous text-to-video generative models, our work focuses on diffusion-based T2V models due to their higher visual quality and better latency in comparison to autoregressive models [19].
-In our work, we argue that the text conditioning mechanism is unified across diffusion T2V models, and thus TALC framework should generalize to various model architectures.
-Specifically, we perform our experiments with ModelScope, a latent diffusion model, and Lumiere, a diffusion model that operates in the raw video space.
-In the future, it will helpful to adopt TALC for multi-scene video generation in novel training and model design paradigms.
-In this work, we evaluate the multi-scene video generation capabilities uptill four scenes.
-This is at par with the number of scenes in prior multi-scene video generation works [18, 23].
-In our experiments, we further show that the performance of the models degrade as we generate more scenes, and hence, we limited ourselves to four scenes.
-A relevant future direction would be to expand the quality of multi-scene video-text data beyond four scenes and finetune T2V models using TALC framework.
-In addition, we perform human annotations using Amazon Mechanical Turkers (AMT) where most of the workers primarily belong to the US and Canada (hourly rate - \$$\$18$ per hour).
-Hence, the human annotations in this work do not represent the diverse demographics around the globe.
-As a result, our human annotations reflect the perceptual biases of the annotators from Western cultures.
-In the future, it will be pertinent to assess the impact of diverse groups on our human evaluations.
-Finally, we acknowledge that the text-to-video generative models can perpetrate societal biases in their generated content [45, 46].
-It is critical that future work quantifies this bias in the generated videos and provides methods for safe deployment of our models.
-Appendix B Detailed Related Work
-Text-to-Video Generative Modeling.
-The field of text-to-video (T2V) synthesis has significantly evolved from its inception with models like VGAN [2] and MoCoGAN [41], leveraging the foundational technologies of GANs [42] and VAEs [43] to produce concise, single-scene videos.
-The narrative depth was further expanded through transformer-based architectures such as CogVideo [47] and VideoGPT [48], enhancing the complexity of video content yet remaining within the confines of single scenes.
-The advent of diffusion models, exemplified by Imagen Video [44], marked a notable advancement in T2V synthesis.
-Despite these strides, the challenge of creating multi-scene videos that reflect the complexity of the physical world [1, 2, 3] remains.
-Our work, TALC, extends the capabilities of T2V models to multi-scene storytelling, filling a crucial gap in the synthesis landscape.
-Image-to-Video Animation.
-The exploration of multi-scene video generation, innovative methods such as Lumiere [11] and Make-a-Video [49] have employed a two-step process, transforming text to images and then animating these images into videos.
-While these approaches have advanced visual quality, they often fall short in weaving seamless multi-scene narratives.
-This limitation is echoed in the work of Emu Video [7], which underscores the difficulty of achieving narrative coherence across multiple scenes.
-TALC focuses on direct generation of multi-scene narratives from textual prompts aiming for a narrative flow and visual consistency across scenes.
-Multi-Scene Video Generation.
-The pursuit of multi-scene T2V synthesis has been furthered by recent innovations like Phenaki [18] and Stable Video Diffusion [9], which have explored new frontiers in video generation from textual prompts and the scaling of latent diffusion models, respectively.
-Additionally, Dreamix [50] and Pix2Video [51] have broadened the scope of diffusion models, applying them to video editing and animation.
-Despite these advancements, the task of generating videos that convey coherent narratives across multiple scenes remains formidable, highlighted by recent works such as VideoPoet [19], ModelScope [10] and Make-A-Scene [52].
-TALC tackles this task and offers a framework produces videos spanning multiple scenes.
-We also introduce nuanced evaluation approach.
-This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis.
-Appendix C Qualitative comparison between baselines and TALC
-We provide an qualitative example to showcase the benefit of TALC over merging captions and merging videos baselines in Figure 6.
-Figure 6: Examples of multi-scene video generation baselines.
-(a) Generating video on the merged descriptions, leads to a poor text-video alignment.
-(b) Generating videos for the individual text descriptions and concatenate them temporally, leads to a lack of visual consistency.
-(c) Our approach (TALC) enhances the scene-level text-video alignment and maintains visual consistency.
-Appendix D Prompt for Multi-Scene Caption Generation
+Appendix A Related Work
+Text-to-Video Generative Modeling:
+Diffusion models like Imagen Video [21] represent a significant advancement in T2V synthesis, yet generating multi-scene videos that realistically capture the complexity of the physical world [1, 47, 10] remains challenging.
+Recent research has attempted longer video generation, but limitations persist.
+Phenaki [45] targets arbitrary-length videos with a focus on temporal coherence, though its evaluation is constrained by the lack of publicly available code.
+VideoDirectorGPT [30], DirecT2V [23], and Free-bloom [24] employ zero-shot approaches for multi-scene generation, contrasting with our fine-tuned method for enhanced performance.
+Gen-L-Video [48] uses iterative denoising to create consistent videos by aggregating overlapping short clips, but it typically generates videos with a maximum of two scenes, whereas our method supports up to four distinct scenes.
+StreamingT2V [19] employs an auto-regressive, streaming-based approach for continuous video generation.
+While these methods contribute to long video generation, they often struggle with maintaining visual consistency and adherence to multi-scene textual descriptions, which are critical for storytelling.
+Multi-Scene Video Generation:
+Efforts like Phenaki [45] and Stable Video Diffusion [7] push the boundaries of text-driven generation by scaling latent diffusion models.
+Dreamix [34] and Pix2Video [12] utilize diffusion models for video editing and animation, while methods like MCVD [46] and VDT [32] focus on improving temporal consistency in longer videos.
+Despite these advances, generating multi-scene videos that accurately reflect complex narratives with high visual fidelity remains difficult, as shown by ongoing research in projects like VideoPoet [28], ModelScope [49], and Make-A-Scene [17].
+TALC addresses the challenges of multi-scene video generation by enhancing visual consistency and text adherence across scenes.
+Unlike zero-shot methods, TALC leverages fine-tuning on a curated multi-scene video dataset.
+We also propose a comprehensive evaluation protocol, combining human evaluation and automated assessments using GPT-4V, to ensure robust narrative coherence in the generated videos.
+Appendix B Evaluator
+Human Evaluation.
+Here, we use the annotators from Amazon Mechanical Turk (AMT) to provide their judgements for the generated videos.
+Specifically, we choose the annotators that pass a preliminary qualification exam.
+Subsequently, they assess the multi-scene generated videos along the dimensions of entity and background consistency, text adherence, and visual quality.
+For each metric, the multimodal model assigns one of three possible response $\{yes=1,partial=0.5,no=0\}$.
+For instance, $yes$ for the entity consistency metric implies that the video frames sampled from the generated video have consistent appearance of the entity described in the multi-scene script.
+We present the screenshot of the UI in Appendix §H.
+Automatic Evaluation.
+Here, we utilize the capability of a large multimodal model, GPT-4-Vision [35], to reason over multiple image sequences.
+First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for two-scene video).
+Then, we prompt the multimodal model with the temporal sequence of video frames from different scenes and the multi-scene text description.
+The model is instructed to judge the generated videos for visual consistency and text adherence, similar to the human evaluation.
+In this work, we do not utilize any existing video-text alignment models [52, 5] for evaluating text adherence as they are trained on single-scene video-text datasets.
+We find that the agreement between the automatic evaluation and human evaluation is $77\%$.
+We present the automatic evaluation prompt in Appendix §G.
+Figure 4: Our approach for generating time-aligned video captions.
+The process begins with PyScene cuts identifying the boundaries of distinct scenes within a video.
+Keyframes are then selected from the median of each scene.
+These frames are processed collectively through the Gemini model to produce multi-image captions that maintain narrative continuity by contextualizing each scene within the video’s overall sequence.
+Appendix C Multi-scene video data sources
+To construct a multi-scene video-text dataset, we utilize existing dataset that include natural (real) videos and associated high-quality human-written captions that summarize the entire video.
+Specifically, we choose MSR-VTT [53] and VaTeX [51].
+Most of the videos in MSR-VTT are 10-30 seconds long while VaTeX consists 10 seconds long videos.
+In addition, each video in MSR-VTT and VaTeX consists 20 captions and 10 captions, respectively, out of which one is selected at random for multi-scene data generation.
+As described above, a single video is cut into multiple video segments using Pyscene library.
+In our experiments, we retain the first four video segments and discard any additional segments if the library generates more than four.
+Since the accuracy of the multi-scene captioning and the computational demands during finetuning are influenced by the number of scenes, we opt to limit the scene count to four for our experiments.
+However, future work could employ similar methodologies to scale the number of scenes, given more computing power and advanced multi-scene captioning models.
+We provide the data statistics for the final multi-scene data in Appendix §K.
+Appendix D Visual quality of the generated videos.
+Table 4: Human evaluation results on the visual quality of the generated videos from ModelScope.
+We observe that the visual quality of the generated videos are close to each other for the base model.
+However, finetuning the model with merging captions reduces the video quality by a large margin while TALC-finetuned model retains the video quality.
+Method Quality
+FreeNoise 80
+Merge captions 80.5
+Merge videos 86.5
+Finetuning w/ Merge captions 63.4
+TALC 84.5
+Finetuning w/ TALC 83.3
+We compare the visual quality of the generated videos using human evaluation in Table 4.
+We find that the visual quality of videos generated from the base model ranges from $80-86.5$ using the baselines and TALC framework.
+However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points.
+This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video.
+Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining.
+Appendix E Prompt for Multi-Scene Caption Generation
 We present the prompt used to generate multi-scene captions using large multimodal model, Gemini-Pro-Vision, in Figure 5.
@@
 Figure 5: Prompt to generate multi-scene caption using large multimodal models.
-Appendix E Source datasets for multi-scene video-text data creation
-To construct a multi-scene video-text dataset, we utilize existing dataset that include natural (real) videos and associated high-quality human-written captions that summarize the entire video.
-Specifically, we choose MSR-VTT [53] and VaTeX [54].
-Most of the videos in MSR-VTT are 10-30 seconds long while VaTeX consists 10 seconds long videos.
-In addition, each video in MSR-VTT and VaTex consists 20 captions and 10 captions, respectively, out of which one is selected at random for multi-scene data generation.
-As described above, a single video is cut into multiple video segments using Pyscene library.
-In our experiments, we retain the first four video segments and discard any additional segments if the library generates more than four.
-Since the accuracy of the multi-scene captioning and the computational demands during finetuning are influenced by the number of scenes, we opt to limit the scene count to four for our experiments.
-However, future work could employ similar methodologies to scale the number of scenes, given more computing power and advanced multi-scene captioning models.
-We provide the data statistics for the final multi-scene data in Appendix §K.
 Appendix F Task Prompts
@@
 V.C. T.A. Overall V.C. T.A. Overall V.C. T.A. Overall
-Merging captions (Base) 95.9 43.8 69.9 99.5 21.5 60.5 81.9 32.0 56.9
-Merging videos (Base) 69.4 71.5 70.5 71.2 82.3 76.7 57.9 58.7 58.3
-TALC (Base) 93.6 48.3 71.0 99.3 57.3 78.3 81.4 36.1 58.8
-Merging captions (F.T.) 94.2 57.1 75.7 94.9 31.2 63.1 50.8 24.3 37.5
-TALC (F.T.) 93.3 62.6 77.9 98.9 76.3 87.6 79.7 47.9 63.8
+Merge captions 95.9 43.8 69.9 99.5 21.5 60.5 81.9 32.0 56.9
+Merge videos 69.4 71.5 70.5 71.2 82.3 76.7 57.9 58.7 58.3
+F.T. w/ merge captions 94.2 57.1 75.7 94.9 31.2 63.1 50.8 24.3 37.5
+TALC 93.6 48.3 71.0 99.3 57.3 78.3 81.4 36.1 58.8
+F.T. w/ TALC 93.3 62.6 77.9 98.9 76.3 87.6 79.7 47.9 63.8
 I.2 Number of Scenes
@@
 V.C. T.A. Overall V.C. T.A. Overall V.C. T.A. Overall
-Merging captions (Base) 93.2 34.4 63.8 92.5 33.0 62.7 87.4 29.9 58.6
-Merging videos (Base) 66.7 69.9 68.3 65.2 71.9 68.6 63.5 70.7 67.1
-TALC (Base) 92.6 54.4 73.5 89.8 48.0 68.9 87.3 39.3 63.3
-Merging captions (F.T.) 87.7 45.7 66.7 83.2 39.5 61.4 60.0 27.4 43.7
-TALC (F.T.) 88.5 66.6 77.5 90.7 64.1 77.4 87.8 56.1 71.9
+Merge captions 93.2 34.4 63.8 92.5 33.0 62.7 87.4 29.9 58.6
+Merge videos 66.7 69.9 68.3 65.2 71.9 68.6 63.5 70.7 67.1
+F.T. w/ merge captions 87.7 45.7 66.7 83.2 39.5 61.4 60.0 27.4 43.7
+TALC 92.6 54.4 73.5 89.8 48.0 68.9 87.3 39.3 63.3
+F.T. w/ TALC 88.5 66.6 77.5 90.7 64.1 77.4 87.8 56.1 71.9
 Appendix J Inference Details
@@
 Noise scheduler DPMSolverMultiStepScheduler
+add FreeNoise details too
 Appendix K Multi-Scene Data Statistics
@@
 prediction type epsilon
-Appendix M Visual quality of the generated videos.
-Table 7: Human evaluation results on the visual quality of the generated videos from ModelScope.
-We observe that the visual quality of the generated videos are close to each other for the base model.
-However, finetuning the model with merging captions reduces the video quality by a large margin while TALC-finetuned model retains the video quality.
-Method Quality
-Merging Captions (Base) 80.5
-Merging Videos (Base) 86.5
-TALC (Base) 84.5
-Merging Captions (F.T.) 63.4
-TALC (F.T.) 83.3
-We compare the visual quality of the generated videos using human evaluation in Table 7.
-We find that the visual quality of videos generated from the base model ranges from $80.5-86.5$ using the baselines and TALC framework.
-However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points.
-This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video.
-Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining.
-Appendix N Single-Scene Video Generation
+Appendix M Single-Scene Video Generation
+Figure 10: Examples of videos generated by TALC with Lumiere, FreeNoise and Gen-L-Video.
+The videos generated using TALC with Lumiere exhibited smooth motion, seamless transitions between text prompts, and high overall quality.
+In contrast, videos created with FreeNoise were of lower quality, with notably less motion.
+For example, in the second video, the background remained static except for darkening, and the jeep did not appear to be in motion.
+Videos generated with Gen-L-Video demonstrated less smooth transitions between prompts and, at times, showed inconsistencies in object representation, as seen in the first video where the man’s appearance changed repeatedly.
+Figure 11: Qualitative examples of videos generated by TALC with ModelScope.
+Videos generated with TALC allow for the creation of videos where a single object performs multiple actions while maintaining visual consistency.
+Furthermore, TALC facilitates the generation of videos with diverse scenes, as demonstrated in the final example.
 To ascertain that our model’s new multi-scene generation function does not detract from its single-scene generation performance, we conducted a series of evaluations using the VBench framework [25].
@@
 Average 0.63 0.59 0.48
-Figure 12: Examples of videos generated by Time-Aligned Captions (TALC)
+Appendix N FreeNoise failure cases
+In this work, we observe that the state-of-the-art method FreeNoise [38] does not perform well on our multi-scene prompts.
+We provide some qualitative examples to show its inability to introduce new content and limited motions as multiple scenes are requested in Figure 12, 13, and 14.
+Figure 12: FreeNoise generation for the prompt: "A wolf howls at the moon in a dense forest.; The wolf prowls stealthily through the underbrush, eyes gleaming.; The wolf catches a scent and follows it eagerly.; The wolf emerges victorious, holding a fresh kill." Explanation: The generated scene shows a wolf standing in a forest, but the key actions described in the prompt are missing.
+The wolf does not prowl, its eyes are not gleaming, it does not follow any scent, and no kill is presented.
+This indicates a significant gap in the model’s ability to convey the sequential and dynamic nature of the narrative.
+Figure 13: FreeNoise generation for the prompt: "A chimpanzee swings effortlessly through the forest canopy.; The chimpanzee gathers fruit, chattering with companions." Explanation: Instead of depicting the chimpanzee swinging through the canopy, the generated scene shows it sitting on a tree branch and then on the ground.
+Furthermore, the scene lacks the critical elements of the prompt, such as the presence of fruits and companions, which are essential for conveying the intended narrative.
+Figure 14: FreeNoise generation for the prompt: "A tiger prowls through dense jungle undergrowth stealthily.; The tiger ambushes a deer, swift and deadly." Explanation: The scene fails to show the tiger ambushing a deer as described.
+Instead, the tiger is depicted standing calmly, and no deer is visible.
+Additionally, a second tiger appears unexpectedly from the first tiger’s body, abruptly changing directions, which introduces further inconsistencies and detracts from the intended narrative.